### v1

# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

Abstract
The Mountain pine beetle (MPB) is a destructive forest pest that undergoes boom-bust cycles every 20-40 years. It also has unusual life-history traits, including a chemosensory affinity for large pine trees, coordinated attacks facilitated by aggregation and anti-aggregation pheromones, and variable dispersal that increases in the absence of trees. Using an empirically calibrated model, we show that MPB’s distinctive life-history characteristics are responsible for its distinctive population dynamics. Specifically, three life-history traits are necessary for episodic boom-bust dynamics: density-dependent dispersal, where beetles disperse once most of the large trees have been killed; an Allee effect, which requires a threshold number of beetles to overcome tree defenses; and a short generation time, notably in comparison with their pine tree hosts. In addition to explaining the qualitative behavior of MPB dynamics, these three ingredients explain residual tree density, the duration of outbreaks, and the average waiting time between outbreaks. The peak number of beetles, believed to be the primary factor driving range expansion, is mostly a consequence of MPB’s constitutive high fecundity. However, two life history traits — MPB’s size-dependent fecundity and preference for large trees — are responsible for 25% of the peak number of beetles. By identifying patterns across the extensive MPB literature, and by integrating data across multiple sources, we develop a highly accurate description of MPB’s density-dependent interactions with lodgepole pine. We conclude with a simplified mechanistic model, distilled to two difference equations, effectively capturing the essence of MPB outbreak dynamics.
Keywords: mountain pine beetle, outbreak, excitable, density dependence
1 Introduction
Over the last 25 years, the mountain pine beetle (MPB; Dendroctonus ponderosae) has killed over 140,000 $\text{km}^{2}$ of pine forests (Fig. 1). Populations of MPB grew synchronously, from Arizona to British Columbia, coalescing to form the largest insect outbreak ever recorded (Taylor et al.,, 2006). By comparison, the previous continent-scale outbreak of MPB in the 1980s appears minuscule. On the local scale (around $1\text{km}^{2}$), MPB populations grow geometrically, plateau in abundance for 1 to 2 years, and then rapidly decline (Fig. D.3).
Figure 1: Time series of MPB damage, as measured by contiguous area of tree crown mortality. To estimate contiguous tree crown mortality, we took fixed-wing Aerial Overview Survey data from both the United States and Canada, used histogram matching to harmonize across different data modalities, and then using the “FHP method” of Egan et al., (2020). The All regions category includes British Columbia (B.C.), Alberta, and Regions 1–6 as defined by the United States Forest Service; 1997 is the first year that all regions were surveyed. The similarity of the All regions and B.C. time series reveals that most of the damage was concentrated in British Columbia.
Figure 2: Progression of the 2000s MPB outbreak, also known as the hyperepidemic. The data comes from fixed-wing Aerial Overview Surveys (United States and British Columbia) and Heli-GPS surveys (Alberta).
Why are mountain pine beetles like this? Why are they so destructive, why are outbreaks widespread, and why are outbreaks episodic with rapid booms and busts? There are several ways to answer these questions, and the relevance of each answer will depend on the underlying point of comparison. For instance, if we ask why the 2000s outbreak was large, relative to past outbreaks, then climate change provides an answer (Carroll et al., 2006a, ; Alfaro et al.,, 2009; Creeden et al.,, 2014). Widespread drought conditions caused synchronized increases in tree susceptibility. At the same time, warmer winter temperatures reduced overwintering mortality in beetle larvae. Fire suppression has also played a role, by enabling the buildup of large-diameter pines (Taylor et al.,, 2006). Alternatively, we might wonder why MPB is so destructive, relative to other insect pests. Most bark beetles are secondary bark beetles, meaning that they infest downed, diseased, or otherwise weakened trees. By contrast, MPB is a primary bark beetle, which means that it attacks and kills healthy host trees, thus accessing all the resources of healthy phloem. It is notable that of the four main bark beetles in British Columbia, the two most destructive (mountain pine beetle and the western balsam bark beetle) are obligate primary beetles, whereas the two least destructive (spruce beetle and Douglas-fir beetle) usually prefer downed trees. A comparative analysis of bark beetles found that the occurrence of outbreaks mostly depends on the availability of susceptible host trees, not life-history traits (Koricheva et al.,, 2012, though the primary/secondary bark beetle distinction was not included in the analysis). Anecdotally, this conclusion is supported by the fact that MPB’s primary host, lodgepole pine, is the most abundant tree in British Columbia, Finally, we might ask why MPB was so much more destructive in British Columbia, relative to the western United States (see Figs.2 & 2). Although this spatial pattern is not well-researched, we suspect that it involves the high density and spatial continuity of lodgepole pine in the Chilcotin Plateau –– the epicenter of the last two outbreaks in British Columbia.
In this paper, we explore how the distinctive life history of the mountain pine beetle affects its population dynamics. We ask: Why do mountain pine beetles experience episodic boom-bust dynamics, relative to an archetypical predator-prey system in which limit cycles or equilibria often appear (e.g., Rosenzweig and MacArthur,, 1963)? To answer this question we 1) build and validate a realistic model of MPB dynamics, and 2) use the model to perform a causal analysis (Fig. 3). The causal analysis involves counterfactual simulations of the model where the traits are turned on/off, followed by a comparison of simulated dynamics and real outbreaks.
Figure 3: Graphical representation of our inferential approach. First, we develop and validate a realistic model of MPB dynamics. Second, we perform a causal analysis in which we turn life-history traits on or off. In the example of the causal analysis, the red crosses respectively indicate that life-history trait # 1 has been off, and that the simulated system does not match reality with respect to Metric #2 (e.g., outbreak periodicity). The dashed red arrow indicates that life-history trait #1 is responsible for realistic values of Metric #2.
1.1 The biology of mountain pine beetle and lodgepole pine
For additional context, we provide a brief natural history of the MPB/lodgepole pine system. More detailed examinations are provided by Safranyik and Carroll, (2006), Raffa, (2001), and Lotan et al., (1983). The univoltine life cycle of MPB begins with adult beetles emerging from their natal trees in July or August. After emergence these beetles fly in search of new host trees, typically targeting large pines. Suitable host species include lodgepole, ponderosa, limber, whitebark, and western white pines. MPB has infested Jack pine and Rocky Mountain bristlecone pine, but the general suitability of these species is a matter of ongoing research (Bentz,, 2019; Bentz et al.,, 2022; Bleiker et al.,, 2023). Individual beetles may fly around for several days, but few beetles survive more than 5–10 days outside of the tree (Reid and Reid,, 2008; Fuchs and Borden,, 1985). Emergence from natal trees is synchronous (due to temperature-dependent development thresholds; Powell and Bentz,, 2009; Logan and Powell,, 2001), but the synchronicity is not perfect, and so the flight period of the entire population may last up to 30 days.
Because individual beetles cannot be tracked, the dispersal phase of the life cycle remains somewhat enigmatic. However, available evidence points to three modes of dispersal:short-, medium-, and long-range dispersal. First, a large majority of beetles disperse less than 30-50 meters under the canopy (Safranyik et al.,, 1992; Robertson et al.,, 2007; for the Douglas-fir beetle Dodds and Ross,, 2002). Second, a minority of beetles disperse up to 5 km (Robertson et al.,, 2009; Powell and Bentz,, 2014; Howe et al.,, 2021; Koch et al.,, 2021), performing pioneer attacks and aggregating to form semi-regular patterns of “spot infestations” (clusters of infested trees) across the landscape (Strohm et al.,, 2013). Finally, 0–2.5% of beetles fly above the canopy (Robertson et al.,, 2009; Safranyik et al.,, 1992), though experts have speculated that the percentage may be higher in stands where most suitable host trees have already been killed (Bleiker,, 2019). These above-canopy beetles may get swept away and carried up to 300 km by upper atmospheric winds (Hiratsuka et al.,, 1982; Cerezke,, 1989; Jackson et al.,, 2008).
Upon locating a suitable host, MPB females initiate attacks. The trees fight back by exuding resin, which serves to entrap or pitch out (i.e. expel) attacking MPB. Female beetles release aggregation pheromones that attract additional beetles. This strategy, employed by many bark beetle species, is called a mass attack. Depending on the beetle population density, tree desirability, and the tree’s defensive capability, successful mass attacks may last from a couple hours to a week, continuing until the tree’s resin supply is depleted. Subsequently, females bore into the tree and begin excavating egg galleries. Mating occurs within the tree, and males will either seek out additional females or “settle down” to help with gallery defense.
Eggs hatch within one to two weeks, and the newly hatched larvae immediately start feeding on the tree’s phloem. Tree death is hastened by MPB’s fungal symbionts, blue stain fungi spp., which colonize the tree’s transport vessels. The beetles carry fungal spores in specialized mouthparts, thereby spreading and inoculating the fungi. In return, the fungi kill the host trees, thus preventing MPB larvae from being killed by future resinosis, were the tree to recover from the initial mass attack (Safranyik et al.,, 1975). Spores that line pupal chambers also serve as a source of nutrition for pre-emergence adult beetles (Reid,, 1958)
In late fall, the larvae enter a dormant state. Despite the accumulation of antifreeze chemicals in their tissue, late-instar larvae face appreciable mortality when temperatures are less than -30°C, and near-100% mortality when temperatures are less than -40°C (Wygant,, 1940; Safranyik and Linton,, 1991; Logan et al.,, 1995). Less extreme temperatures can kill the less resistant life-stages (i.e., early-instar larvae and pupae), particularly during cold snap in the fall or spring (Régnière and Bentz,, 2007; Rosenberger et al.,, 2017). Come spring, surviving larvae resume feeding and pupate in the late spring or early summer. Adults emerge in the summer, completing the univoltine life cycle. It is around this time, approximately one year after the initial attack, that the foliage of all infested trees has turned a telltale rust-red color. Within 2–4 years after the initial attack, most needles will have dropped from the tree (Klutsch et al.,, 2009; Audley et al.,, 2021).
Lodgepole pine and ponderosa pine are the predominant hosts of MPB, simply because they are the most abundant pine species in MPB’s geographic range and elevational niche. Lodgepole pine dominates in British Columbia, western Alberta, and the American Rocky Mountain states, whereas ponderosa pine is more abundant in the American southwest and the Sierra mountains (Beaudoin et al.,, 2014; Ellenwood et al.,, 2015). Both species are shade-intolerant and fire-resistant. They tend to be seral (i.e., persisting as a non-climax species in a fire cycle), taking advantage of high-light conditions following a stand-leveling fire (Fryer,, 2018; Anderson,, 2003; Lotan et al.,, 1985). Because MPB rarely kills all of the mature trees in a stand, and because pines produce many cones, no pine species faces recruitment limitation during recovery from MPB outbreaks. This is doubly true for lodgepole pine, which has closed cones with seeds that remain viable for decades, effectively creating a seed bank (Teste et al., 2011a, ; Teste et al., 2011b, ). Due to its abundant seeds and its ability to thrive in high-light conditions, lodgepole pine can take advantage of the light gaps created in the years following a mountain pine beetle attack, resulting in the recruitment of seedlings into approximately even-aged cohorts (Axelson et al.,, 2010)
During epidemics, MPB seeks out large healthy trees with abundant phloem resources. However, in their low-density endemic state, there are not enough beetles to mass attack a single healthy tree, and thus there is a collective behavioral switch where beetles seek out trees that are weakened or stressed, typically due to root disease, wind damage, or lightning (Tkacz and Schmitz,, 1986; Carroll et al., 2006b, ). Endemic populations tend to remain at low density, primarily due to resource limitation, but also because of parasitoids, woodpeckers, and competition with secondary bark beetle species (Amman,, 1983; Safranyik and Carroll,, 2006). Endemic populations of mountain pine beetle (MPB) can transition into an incipient-epidemic state when they successfully breed within trees that have thick phloem. This scenario can occur if a healthy tree experiences acute stress (such as wind damage or lightning) right before the MPB flight period, such that secondary bark beetles have not yet extensively colonized the tree (Bleiker et al.,, 2014). Such stressors can trigger local outbreaks (on the scale of square kilometers). However, widespread outbreaks are usually preceded by widespread drought conditions, which synchronously weaken the resin defenses of many trees (Alfaro et al.,, 2009; Creeden et al.,, 2014).
1.2 Mountain pine beetle life history traits
A key element of this paper is a nonlinear dynamical model that captures key MPB life history traits. In this section, we review the traits with a view towards qualitative features that should be emphasized in the model. Here, life history traits refers to the nonstandard aspects of MPB’s host-parasitoid interaction with lodgepole pine, including the ways in which MPB “decides” about what trees to attack, and how MPB responds to variation in the density of host trees and conspecific beetles. Our list of life history traits is not exhaustive. Previous research has identified some traits that do not show up as a clear pattern in our data, or traits that we cannot feasibly incorporate into our models. These are discussed further in Appendix B.
It is worth keeping in mind that what counts as a life history trait is somewhat a matter of perspective. As we will discuss, both MPB’s stand-dependent dispersal and large-tree attack preference could potentially stem from its inability to chemically sense small trees; are these two separate life-history traits, or one combined trait? We take a pragmatic approach, favoring traits that are conceptually separable and traits that can be independently manipulated in our mechanistic model.
1.2.1 Allee effect
Mountain pine beetle (MPB) is subject to an Allee effect, a phenomenon where individual fitness increases with population density. The Allee effect arises because a threshold number of beetles must attack a single tree to successfully kill it, but this threshold is challenging to reach in low-density populations. Another common mechanism of an Allee effect, poor mate finding in small populations, has been ruled out by previous research (Amman,, 1980; Bleiker et al.,, 2014).
The Allee effect is mitigated by MPB’s propensity for aggregation. In general, aggregation counteracts the negative effects of positive density-dependence, just as dispersal counteracts the negative effects of negative density dependence (Crowley,, 1981; Briggs and Hoopes,, 2004). Females release aggregation pheromones and attacked trees emit terpene volatiles, which together attract additional beetles. However, this chemical signal is dampened by unsuitable host-tree chemistry (Hynum and Berryman,, 1980; Boone et al.,, 2011), unsuitable natal-tree chemistry (which negatively affects aggregation pheromone production; Chiu et al.,, 2018), and a limited detection limit for aggregation pheromones, which is around 100-1000 meters (Barclay et al.,, 1998, Fig. 2; Strohm et al.,, 2013, Fig. 4; Mitchell and Preisler,, 1991, Fig. 4–5). Thus, the Allee effect persists due to various physical and chemical factors that constrain the strength of aggregation.
It has long been recognized that the size disparity between beetles and trees means that many beetles must attack a single tree, and therefore, that outbreaks require a threshold or critical density of beetles (Thalenhorst,, 1958; Safranyik et al.,, 1975; Berryman,, 1978). However, explicit evidence for the Allee effect in MPB is not always forthcoming. Studies have detected an Allee effect when focusing on tree-level attack density (e.g., beetle entry holes per unit tree surface area). An Allee effect is implied by a positive relationship between attack density and the probability of tree death (Waring and Pitman,, 1983, 1985; Peterman,, 1974), a positive relationship between attack density and the probability of reproduction (Raffa and Berryman,, 1983, Fig. 5), or a positive relationship between the number of attacked trees and the probability of tree death (Boone et al.,, 2011). On the other hand, studies do not detect an Allee effect when they focus solely on the density of infested trees, treating this quantity as a proxy for beetle population size (MacQuarrie and Cooke,, 2011; Tunnock et al.,, 1970; Trzcinski and Reid,, 2009).
There are at least three explanations for why the Allee effect is not evident in some datasets. 1) Within-tree beetle density is positively related to the probability of tree death and the likelihood of beetle reproduction, but the counter-veiling impact of negative density dependence cancels out this signal. Beetle density at the tree level is negatively related to per capita oviposition and per capita survival (Raffa and Berryman,, 1983, Fig. 10). Further, at the stand level, high beetle density suggests that many trees have been killed and that predominantly low-quality trees remain. Thus, the demographic Allee effect at the stand level is weaker than the component Allee effect at the tree level. 2) While aggregation counteracts the negative effects of positive density-dependence, it also makes it difficult to observe a low-density but non-zero population at the stand level. To see this, consider the hypothetical scenario where aggregation was perfect (i.e., beetles can travel and communicate perfectly over long distances). In this scenario, you would never observe a tree with a low attack density, so you could never uncover a positive relationship between attack density and tree death. Indeed, tree-level studies fail to uncover an Allee effect when they only examine successfully attacked trees (Reid,, 1963; Berryman,, 1974, Fig. 2c; Berryman,, 1976, Fig. 2). We don’t live in the perfect aggregation scenario — in reality, we observe some trees with a sub-threshold attack density — but perhaps aggregation is strong enough that we rarely observe stands in which the overall beetle density is below the theoretical stand-level allee threshold. 3) The majority of beetles are thought to disperse less than 30 meters, so by calculating the density of infested trees over acres or hectares, the aforementioned studies (MacQuarrie and Cooke,, 2011; Tunnock et al.,, 1970; Trzcinski and Reid,, 2009) are calculating a spatial average over a number of weakly coupled sub-populations. Averaging effectively reduces the variation in beetle density, and in turn, makes it more difficult to detect an Allee effect. This statistical stabilization hypothesis has been used to explain weak population regulation across many species (De Roos et al.,, 1991; Ray and Hastings,, 1996; Thorson et al.,, 2015).
1.2.2 Stand-Dependent Dispersal
MPB predominantly disperses from stands only after most large-diameter trees are killed. The vast majority of these emigrants die on their journey, either due to an Allee effect or a low energetic condition at their destination. Since stand-dependent dispersal and the associated mortality are somewhat obscure features of MPB dynamics, we will take a moment to discuss supporting evidence.
There are several potential explanations for stand-dependent dispersal. One hypothesis suggests that MPB is simply unable to aggregate effectively on small and medium-sized trees, putatively due to chemosensory constraints. Field data support this — MPB primarily attacks small trees when they are adjacent to large trees (Mitchell and Preisler,, 1991; Preisler and Mitchell,, 1993), analogously to how MPB attacks nonhost trees only when they are adjacent to heavily-infested host trees (Amman et al.,, 1985; Huber et al.,, 2009). Further, studies have shown large trees have higher monoterpene concentrations, which may be the main determinant of whether pioneer female beetles initiate an attack (Boone et al.,, 2011; Seybold et al.,, 2006). A secondary explanation considers how beetle infestations alter the forest microclimate (Bartos,, 1989; Amman and Logan,, 1998). As stands are thinned, increased wind disrupts MPB aggregation pheromone plumes, and increased light may induce MPB to fly. A final hypothesis claims that beetle stand-dependent dispersal is a behavioral adaptation that maximizes inclusive fitness. Large trees have thicker phloem leading to high per capita fecundity, so lineages that seek out large trees may have higher geometric fitness than their less-risky competitors, despite the high probability of dispersal mortality. To our knowledge, this explanation has not yet been explored by the literature.
We propose that stand-dependent dispersal in beetles is best explained by the chemical-sensing hypothesis, i.e., beetles are less effective at mass attacking small trees. This tentative conclusion was reached by a process of elimination. The problem with the microclimate hypothesis is that there is a time lag between a beetle infestation and biotic forest thinning. Most MPB-killed trees don’t fall down for 5-20 years (Lewis and Hartley,, 2006; Hansen,, 2014), and most don’t lose their needles for 2–7 years (Klutsch et al.,, 2009; Audley et al.,, 2021), but the populations crashes occur 1–2 years after infestations have peaked (Fig. D.3). The inclusive fitness hypothesis is unlikely because disperses face significant mortality, and because beetles have positive expected fitness when they do successfully kill medium-sized trees (Fig. 4).
Even though we cannot track individual beetle dispersal, there is robust evidence for stand-dependent dispersal. In addition to the previously discussed evidence, pheromone-trap experiments show substantially higher numbers of trapped beetles in thinned vs. unthinned stands (Bartos,, 1989; Schmitz et al.,, 1989). Sallé and Raffa, (2007) have shown that beetles with low body weight had proportionally more lipids. Since beetles are more reactive to aggregation pheromones once lipids are depleted, this explicates a physiological mechanism by which beetles can escape depleted stands. Several modelling studies explicitly infer substantial stand-dependent dispersal(Powell and Bentz,, 2014; Powell et al.,, 2018; Howe et al.,, 2021). Other, more complicated models implicitly assume stand-dependent dispersal, and then show that the model generates realistic behavior (e.g., Powell et al.,, 2018; Bone and Altaweel,, 2014; Kautz et al.,, 2016).
The most compelling evidence for stand-dependent dispersal is the fact that outbreaks invariably end before the vast majority of medium-sized trees are killed (Fig. 5), even though beetles have positive fitness on medium-sized trees (Fig. 4). Previous research (Goodsman et al.,, 2018) has explained this pattern —– where an outbreak ends sooner than expected —– as a result of overcompensatory density dependence, a phenomenon where increased population size now leads to a decreased population size in the next generation. However, there is not strong evidence for overcompensatory density dependence in MPB. Attacking beetles pack themselves into trees at a density that optimizes individual fitness (Raffa and Berryman,, 1983). When attack densities do exceed the theoretical optimum, there is no clear change in emergence densities (Fig. D.2; Berryman,, 1974, Fig. 2c; Berryman,, 1976, Fig. 2, Raffa and Berryman,, 1983, Fig. 6 shows a slight increase) or population dynamics (Appendix A.4).
Figure 4: Mountain pine beetle fitness as a function of diameter at breast height (DBH), in inches. Ecological fitness is estimated with (0.3)×(0.3$\times$( 0.3 ) × females out of the tree / females into the tree, where the factor $0.3$ is the approximately the median dispersal-phase survival in the dataset from Klein et al., (See Fig. D.1). Depending on the study, gallery starts or entry holes serve as estimates for females into the tree (note that only females initiate galleries). (2/3)×(3/2)×(2/3$\times$(3/2$\times$( 2 / 3 ) × ( 3 / 2 ) × Emergence holes, or $(2/3)$ of the overwintered brood serve as estimates of females out of the tree. Note that (2/3) is the typical proportion of females (Reid,, 1962; Cole et al.,, 1976), and that (3/2) is the average number of beetles emerging from each hole (Safranyik and Linton,, 1985; Peterman,, 1974, Fig. 13). Data comes from Goodsman et al., (2018); Cudmore et al., (2009); Reid, (1963); Safranyik and Carroll, (2006). The means and standard errors were only computed for bins with at least 5 observations.
Figure 5: The proportion of remaining live trees as a function of time, stratified by diameter class (diameter at breast height, DBH). Data comes from Stipe, (1974, 1976); Gibson and Bennett, (1978); Gibson et al., (1979); Klein et al., (1978); Preisler and Mitchell, (1993); Nelson, (2009).
There is also evidence that stand-dependent dispersal mostly results in mortality, although this evidence is indirect. First, it is our observation (based on Aerial Overview Survey data) that outbreaks do not last substantially longer in areas with high contiguous pine biomass. If dispersing beetles were not dying in such areas, emigrating beetles would be replaced by immigrating beetles, and thus outbreaks would be prolonged. Second, mortality is already substantial for beetles traveling short distances: baseline dispersal mortality is estimated to be around 40–80% (Safranyik,, 1989; Latty,, 2007; Pope et al.,, 1980); Fig. D.1). Third, dispersal puts beetles in a low energetic condition, leading to decreased aggregation pheromone production, and making them less likely to attack trees (Jones,, 2019; Jones et al.,, 2020). Fourth, lab experiments indicate that most beetles only survive several days outside of a tree (Fuchs and Borden,, 1985; Reid and Reid,, 2008).
1.2.3 Size-dependent attack preference
Large trees offer several benefits to MPB, so it is no surprise that MPB has a strong preference for large-diameter trees in the landing, pioneer attack, and mass attack phases of its life cycle. In experiments, beetles prefer landing on large, vertically-oriented black silhouettes (Shepherd,, 1966; Billings et al.,, 1976), indicating that visual cues play a major role in initial, pre-aggregation landings. Chemo-sensory cues play a secondary role in the initial landing phase (Billings et al.,, 1976; Hynum and Berryman,, 1980; Moeck and Simmons,, 1991). After landing, a female pioneer beetle decides whether to attack based on a combination of physical and gustatory cues associated with tree diameter, such as terpene concentrations (Raffa and Berryman,, 1982) and bark roughness (Reid,, 1963). Once mass attacks have been initiated, both males and females are attracted to large trees, due to a synergist attraction to aggregation pheromones and volatilized terpenes, both of which scale with tree size.
MPB’s large-tree attack preference is extremely well-established. The probability of tree death increases logistically with tree diameter (Buonanduci et al.,, 2023; Rodman et al.,, 2022; Morris et al.,, 2022; Negrón,, 2020; Negrón et al.,, 2017; Björklund and Lindgren,, 2009 and 15 sources therein, see Table 1), and stand susceptibility indexes invariably use variables that correlate with tree diameter: tree age, pine volume, or basal area (Shore et al.,, 1992; Chojnacky et al.,, 2000; Randall et al.,, 2011; Bentz et al.,, 1993, and sources therein). Depending on site-specific factors, the probability of tree mortality is near-zero for 4–8-inch trees and near-100% for 12–16-inch trees.
1.2.4 Size-dependent fecundity
MPB has higher effective per capita fecundity in larger trees. Large trees have thick phloem, which is the primary food source of larvae and teneral adults. Thin phloem can slow larval development, which can cause mortality if a cold snap occurs before the cold-resistant third & fourth-instar stages are attained; if larvae do manage to fully develop within a small tree, they may have small body sizes and low energetic condition upon emergence (Amman,, 1983). Additionally, the thick bark of large trees protects against cold winter temperatures.
From a more abstract perspective, there are two ways in which large trees lead to more beetles: 1) they offer higher emergence densities (per unit of susceptible tree surface area); 2) they have lots of susceptible surface area. We use the term effective fecundity specifically to the former, diameter-dependent emergence density; we use “effective” because tree diameter doesn’t affect brood density per se, but rather brood density after overwintering but before the dispersal phase. Emergence density has a piecewise-linear relationship with tree diameter (Reid,, 1963, Table 1; Safranyik,, 1968, Fig. 25; Cole,, 1969, Fig. 5; Cudmore et al.,, 2009, Fig. 6). Small trees produce next-to-zero beetles, whereas large trees produce thousands of beetles per square meter. Susceptible tree surface area also increases linearly with tree diameter, implying that total beetle brood size increases quadratically with tree diameter.
1.2.5 Time-scale separation
Trees live longer than beetles. Although this is evident, it sets the MPB-lodgepole system apart from most predator-prey and host-parasitoid systems. Typically, predators live longer than prey, and parasitoids have lifespans that are similar to their hosts’ lifespans.
2 Methods
2.1 Overview
To understand the effects of MPB life-history traits on MPB population dynamics, we will build a mechanistic model of MPB dynamics (Fig. 6), then simulate the model under counterfactual scenarios where various life-history traits have been turned on or off. If a phenomenon (e.g., episodic outbreaks) disappears when a certain life-history trait is turned off, we can say that the life-history trait is a necessary condition for the phenomenon. If two traits being turned off makes the difference, then the two traits are jointly necessary. Of course, our inferences are conditional on nature following the dynamics described by the model.
Figure 6: Graphical representation of key model processes.
Realism is the key feature of our model — we intend to emulate the actual ecological system so that we can conduct virtual experiments. Many models of MPB have been published, but many are too simple for our purposes, having favored mathematical tractability over realism. Other models are too complex, containing functional forms that cannot be reasonably parameterized given available data. For example, while we understand that beetles are attracted to a suite of pheromones, explicitly modelling the production and diffusion of these pheromones — which depend on the stand structure, transient wind conditions, and the natal hosts of attacking female beetles — is a herculean task. Fitting such a model requires that we fix a large proportion of parameters (e.g., Biesinger et al.,, 2000, Table 2), yet we lack the appropriate kind of data to select these parameter values with confidence. More generally, the pursuit of realism in our model is tempered by the inherent complexity of the MPB system and data limitations.
In this section, our overarching principle is that all model structures be thoroughly justified via clear patterns in the data and/or a preponderance of evidence from previous research. Beyond the justifications outlined in the model description (below), additional support is provided in Appendix B. Our model, which examines the interactions between MPB and lodgepole pine over a one square mile area, is validated and parameterized using data from this spatial scale. Given that beetles primarily disperse short distances (refer to Section 1.1), inter-patch dispersal has minimal impact on local populations and is not explicitly modeled. Additionally, some processes related to within-stand dispersal are implicitly modelled. For example, the aforementioned diffusion of pheromones is subsumed under a linear relationship between beetle preference and tree diameter, since larger trees produce larger plumes of attractive terpenoids.
2.2 Data
The model was predominantly validated and parameterized using data from Boone et al., (2011) and Klein et al., (1978). All data was extracted from figures and tables using the GraphGrabber software and Microsoft Excel’s digitizing features. Specific variables are describe in Table A.1. The Boone et al., dataset contains the number of mass attacked trees and the number of successfully mass attacked trees, across a range of attack years, stands, and attack densities. This data justifies a functional form for the Allee effect. The second dataset, courtesy of Klein et al., (Fig. 7), is one of the only datasets that contain beetle attacks, beetle emergence, and tree deaths, all stratified by year and tree diameter class. Our statistical model of MPB outbreak dynamics is primarily calibrated with the Klein et al., dataset, with the notable exception being that a Bayesian prior for the Allee effect parameter is derived from the Boone et al., data (more details in Section 2.3).
Boone et al.,’s study was conducted from 2000–2005 in the mature lodgepole pine stands of southern British Columbia. Six stands were selected for the study, with an average area of 15.3 hectares per stand, and an average of 1225 canopy trees per hectare. Each year, until the beetle population waned and attained its endemic condition, every tree in each stand was inspected to determine the number of attacked trees and the success or failure of these attacks.
Klein et al.,’s study was conducted from 1965–1972 in the southwest corner of Yellowstone National Park in Idaho. Data was collected and aggregated across 85 permanent plots that spanned a square mile. Trees were grouped into six 2-inch diameter classes, ranging from 8 to $\geq$ 16 inches. For each of the 106 sampled trees, paired 6x6-inch sample squares were marked on opposite sides of the bole at regular height intervals. Within these squares, emergence holes and successful attacks (operationalized as egg-gallery starts) were measured. Some of this data was originally presented by Parker, (1973), but a more thorough reporting is given by Klein et al., (1978).
Figure 7: Time series of emergence holes and gallery starts in the Klein et al., dataset. Emergence and subsequent attacks are assumed to take place in July and August respectively. There are always far fewer gallery starts than emergence holes, suggesting high levels of dispersal mortality.
2.3 Simulation model description
2.3.1 Outbreak dynamics
Throughout the model description, we will refer to the surface area over which beetles may attack as the susceptible surface area; note that beetles do not attack the upper part of the bole due to thin phloem. To convert between tree diameter and susceptible surface area, we use a simple linear model,
$SA=-3.792+0.821\times DBH,$ (1)
where $DBH$ is the diameter at breast height in inches, and $SA$ is the susceptible surface area in square meters. Eq. 1 was first published by Burnell, (1977), who claims an excellent fit ($R^{2}=0.94$) and attributes the parameterization to William Klein.
We start with beetle density in year $y$, denoted $B_{y}$ (units: females per acre). We do not track males, since previous research has shown that nearly all females are mated (Reid,, 1958; Amman,, 1980). Upon emergence, beetles disperse in search of a new host. The probability of survival during this dispersal phase is
$s_{y}=s_{0}\lambda_{y},$ (2)
where $s_{0}$ is the maximum survival probability, and $\lambda_{y}$ is a factor (between 0 and 1) that reduces the maximum survival probability when many trees have been killed, putatively because beetles are dispersing farther and thus suffering from higher dispersal mortality.
The effect of stand-dependent dispersal on the dispersal-phase survival is
$\lambda_{y}=\left(1+\exp\left(-\lambda_{1}\left((\text{relative surface area})_{y}-\lambda_{0}\right)\right)\right)^{-1}.$ (3)
The right-hand side is simply a logistically-increasing function of the independent variable, $(\text{relative surface area})_{y}$, which is defined as the susceptible surface area of all trees with $DBH\geq 8\,\text{inches}$ in year $y$, divided by the total susceptible surface area when MPB are absent and the forest is at equilibrium. We use the relative surface area mostly for interpretability but also for computational reasons (the Bayesian model-fitting program Stan prefers all parameters to be on the unit-scale). The parameter $\lambda_{0}$ is the threshold for dispersal (i.e., $\lambda_{y}<0.5$ when $(\text{relative surface area})_{y}<\lambda_{0}$), and $\lambda_{1}$ controls the slope of the logistic function.
We cannot directly justify the functional form of Eq. 3 — say, by plotting total emergence against subsequent total attacks — because Klein et al., only reported attacks on successfully killed trees. However, we note that the logistic function is a standard way to model probabilities, and that the extreme limit of the function (i.e., dispersal mortality is near 100% when there are few trees) is supported by both the Klein et al., dataset (Fig. D.1) and the literature (see Section 1.2.2). In particular, Powell and Bentz, (2014) estimated that mean dispersal distance increased from several meters to several kilometers as tree density decreased.
To obtain the density of attacking beetles, we multiply the emerging females’ density and the dispersal-phase survival probability. Potentially, with probability $P_{I}$, we also add $I$ beetles per acre:
$(\text{{attacks per acre}})_{y}=s_{y}\;B_{y}+I\times\text{Bernoulli}\left(P_{I}\right).$ (4)
The Bernoulli random variable simulates the stochastic endemic $\rightarrow$ incipient-epidemic transition. Because most MPB outbreaks occur during continent-scale epidemics, dispersal is the proximal catalyst for the endemic $\rightarrow$ incipient-epidemic transition on the local scale. But because drought is responsible for region-wide outbreaks, drought is the ultimate catalyst for most local-scale outbreaks. Drought conditions are somewhat unpredictable, but the best predictor of drought in western North America is the Pacific Decadal Oscillation (McCabe et al.,, 2004; Hidalgo,, 2004), which has a quasi-period of 15-30 years (Mantua and Hare,, 2002). The last two drought periods preceded the last two continent-scale outbreaks and were 20 years apart, (Creeden et al.,, 2014, Fig. 5d). Therefore, we set $P_{I}=1/20=0.05$.
To simulate the population size corresponding to the incipient-epidemic phase, we set $I=300$. This number was obtained with a back-of-the-envelope calculation: the median emergence density of females across several studies and Forest Service reports (Other Studies in Fig. 16), multiplied by the susceptible surface area of typical killed tree (10 inches DBH as an input to Eq. 1), multiplied by the median estimate of the dispersal-phase survival probability (using gallery-starts/emergence-holes quotient from the Klein et al., dataset, Fig. D.1), multiplied by 0.81, which is the number of successfully attacked trees per acre that corresponds to an incipient-epidemic state (according to Carroll et al., 2006c, ). The chosen values of $P_{I}$ and $I$ are educated guesses, but we re-ran our analysis with several different values and obtained qualitatively identical results.
After the dispersal phase, the remaining beetles attempt to mass attack pine trees, with the primary determinant of success being the number of nearby beetles. If beetles are unsuccessful, the tree survives and the beetles are assumed to die. This is a simplifying assumption because unsuccessful beetles could attack other trees; however, it is a decent assumption because some beetles will be killed by the first tree, the remainder face increased mortality risk during secondary dispersal, and late-arriving beetles have low fitness (Raffa,, 2001).
[…]
where $a$ is the Allee threshold parameter. More precisely, $a$ is the beetle attacks per acre where the probability of success equals $0.9$. Thus, the number 2.3 in the above equation is obtained by solving $0.9=1-\exp(-x)$ for $x$.
The exponential saturating function in Eq. 5 is justified by the Boone et al., dataset (Fig. 8). A notable discrepancy is that the maximum probability of success reaches (close to) 100% in the model, but not in reality. However, an additional parameter for the maximum probability of successful attacks would be non-separable with the parameter $s_{0}$ (in Eq. 2). Thus, our estimate of $s_{0}$ not only represents maximum survival during the dispersal phase, but also the fact that some mass attacks are unsuccessful despite high beetle density. Additionally, $s_{0}$ captures low-vigor females that successfully attack trees but do not oviposit (5–30% of females could fall into this category; Amman,, 1975, 1980).
Figure 8: The probability of an attacked tree being killed is an increasing function of beetle density. The data is effectively modelled by an exponential saturating function (see Eq. 5, “Probability of success” term). The prediction was generated using posterior means of model parameters. Data from the Boone et al.,.
Next, the successfully attacking beetles distribute themselves between different tree diameter classes. To visualize the extent of MPB’s attack preference, we can plot the total beetle attacks across all trees within a diameter class, as a function of the proportion of tree surface area belonging to that diameter class (Fig. 9). If total beetle attacks are proportional to total tree surface area, then observations should cluster around the one-to-one line. Instead, we see that large trees are favored and small trees are disfavored.
Figure 9: Beetles disproportionately kill large trees. The x-axis is the proportion of susceptible surface area belonging to live trees of a particular diameter class, within a year. The y-axis is the proportion of total attacks received by the same diameter class, within a year. The black line is a LOESS model prediction. The grey dashed line is the one-to-one line, representing the no-mass-attack-preference hypothesis. Data from Klein et al., (1978).
Figure 10: Beetle preference is an increasing function of tree diameter. The data (from Klein et al.,, 1978) is well-modelled by a piecewise linear function (see Eq. 6). There is clear heteroskedasticity in the data, which is likely due to smaller sample sizes within classes of large-diameter trees. The exact number of samples is not available, so we use the within-DBH-class surface area as a proxy for sample size, visualized as scatter-plot point size. The prediction was generated using posterior means of model parameters.
We can operationalize mass attack preference by dividing the proportion of the beetle attacks by the proportion of tree surface area; that is, by dividing the y-coordinate by the x-coordinate in Figure 9. Note that this quantity is equivalent to the relative attack density: the attack density within a particular diameter class, divided by the attack density across all trees (within a year). Mass attack preference has a clear linear relationship with tree diameter (Fig. 10), which justifies the piecewise linear function,
$(\text{preference})_{j}=\begin{cases}\beta_{0}+\beta_{1}\cdot\text{{DBH}}_{j}&\text{if }\text{{DBH}}_{j}>-\beta_{0}/\beta_{1}\\ 0&\text{otherwise},\end{cases}$ (6)
where $\beta_{0}$ and $\beta_{1}$ are the intercept and slope parameters respectively. The subscripts $j$ and $y$ denote the DBH class and the year index respectively. The quantity $-\beta_{0}/\beta_{1}$ is the susceptible diameter threshold, i.e., trees with $\text{DBH}<-\beta_{0}/\beta_{1}$ are not attacked.
[…]
The equation for successful attacks, Eq. 7, shows how beetles are distributed among tree diameter classes, but not how beetles are distributed among individual trees. As it turns out, killed trees receive an approximately invariant number of attacks per unit surface area. That is, attack density does not depend on tree diameter (Fig. 11, Reid,, 1963, Table 1, Peterman,, 1974, Figs. 4–7). There is less literature on the relationship between attack density and beetle population size, but any such relationship is likely weak, given that 1) Negrón, (2018) found no relationship; 2) there is a clear mechanism for limiting attack density (see the following paragraph), which explains 3) that attack density does not depend on tree diameter, and 4) that attack density varies in the range of 50-100 females per square meter, despite beetle population sizes varying by several orders of magnitude. In the Klein et al., dataset, the link between attack density and beetle population size is inconclusive, but this has a negligible effect in any case (Appendix B.2).
Figure 11: Successful attack density (gallery starts per square meter) increases slightly over time, but does not depend on tree diameter at breast height (DBH). Data from Klein et al., (1978).
The mean attack density of 71 females per square meter in the Klein et al., dataset is close to the optimal value of 60 females per square meter, proposed by Raffa and Berryman, (1983). The mechanism for attaining near-optimal attack density is a suite of aggregation and anti-aggregation pheromones, and the relationship between such pheromones and host tree defenses. Beetles are attracted to the combination of terpenoid aggregation pheromones and volatilized terpenes. Because both are derived from tree resin, trees become less attractive as their resin defenses are depleted. Anti-aggregation pheromones, which are only detected by beetles at short distances (around a meter; Bentz,, 1996), redirect beetles to nearby trees, while rivalry stridulations (noises produced by males) help ensure that galleries are evenly spaced. Raffa, (2001) puts forward several hypotheses for the apparent lack of cheaters — beetles who ignore the anti-aggregation pheromones and enter a “full” tree.
Because of the invariant attack density, the number of killed trees per acre in the $j^{\text{th}}$ diameter class, denoted by $(\text{KTPA})_{j,y}$, is given by a linear function
$(\text{KTPA})_{j,y}=\text{min}\left(\gamma\frac{(\text{succ. attacks per acre})_{j,y}}{(\text{surface area per tree})_{j,y}},(\text{TPA})_{j,y}\right),$ (8)
which clearly fits the data (Fig. 12). The slope parameter $\gamma$ has units $m^{2}/\text{attacks}$, so $1/\gamma$ is an estimate of attack density. The maximum number of killed trees is capped by the pre-attack trees per acre, $(\text{TPA})_{j,y}$.
The expected beetles in the following year is obtained by assuming that emergence density (per $\text{m}^{2}$) is a piecewise linear function of tree diameter, and then by summing across all $J$ diameter classes:
$\mathbb{E}\left[B_{y+1}\right]=\sum_{j=1}^{J}\begin{cases}\overbrace{\left(\zeta_{0}+\zeta_{1}\cdot\text{DBH}_{j}\right)}^{\text{Emergence density}}(\overbrace{(\text{KTPA})_{j,y}\cdot(\text{surface area per tree})_{j}}^{\text{surface area of killed trees}}),&\text{if }\text{DBH}_{j}>-\frac{\zeta_{0}}{\zeta_{1}}\\ 0,&\text{otherwise}.\end{cases}$ (9)
Here, $\zeta_{0}$ and $\zeta_{1}$ are the intercept and slope parameters of the emergence $\sim$ DBH regression. The quantity $-\zeta_{0}/\zeta_{1}$ is the beetle-production threshold, i.e., trees with $\text{DBH}<-\zeta_{0}/\zeta_{1}$ will not produce any beetles, even if they are successfully attacked.
Figure 12: The number of killed trees is a linear function of successful beetle attacks (corrected for DBH-specific surface area), suggesting that beetles “fill up” a tree to a universal density, and then move onto the next tree. The prediction was generated using posterior means of model parameters. Data from Klein et al., (1978).
We cannot justify the functional form of Eq. 9 with graphical evidence, since the Klein et al., dataset only contains total emergence holes (not stratified by diameter class). However, the linear nature of the relationship has been well-established (Reid,, 1963, Table 1; Safranyik,, 1968, Fig. 25; Cole,, 1969, Fig. 5; Cudmore et al.,, 2009, Fig. 6). The linear relationship also implicitly accounts for the fact that the male:female sex ratio of emerging beetles increases with tree diameter (Cole et al.,, 1976, Fig. 5). Note that emergence only depends on killed trees, not attacking beetles, because a super-optimal attack density is balanced by a decrease in per capita oviposition or early-instar survival; this is perfectly compensatory density dependence at the tree scale (Fig. D.2; Berryman,, 1974, Fig. 2c; Berryman,, 1976, Fig. 2).
To capture temporal variation in overwintering mortality, we assume that the realized number of beetles is drawn from a zero-truncated normal distribution:
$B_{y+1}\sim\text{Normal}_{\scriptscriptstyle{Trunc.}}\left(\mathbb{E}\left[B_{y+1}\right],\sigma_{\text{env}}\mathbb{E}\left[B_{y+1}\right]\right).$ (10)
The scale of environmental stochasticity is fixed at $\sigma_{\text{env}}=0.2$, which is roughly consistent with the ±$\pm 40\%$ deviations of survival seen in the literature (Safranyik and Vithayasai,, 1971; Reid,, 1963, Tables 4 & 7; Amman,, 1983, Fig. 34; Régnière and Bentz,, 2007, Table 1).
2.3.2 Lodgepole pine dynamics
Each time-step of the simulation model begins and ends with the beetle flight period in July or August. Therefore, the first process affecting the lodgepole pine population is MPB-induced mortality. The density of remaining trees in late summer, denoted $(\text{TPA}^{\prime})_{j,y}$, is
$(\text{TPA}^{\prime})_{j,y}=(\text{TPA})_{j,y}-(\text{KTPA})_{j,y},\quad j=1,\ldots,J.$ (11)
Recall that $(\text{TPA})_{j,y}$ is the initial number of trees per acre, $(\text{KTPA})_{j,y}$ is the number of MPB-killed trees per acre, and $j=1,\ldots,J$ indexes the tree diameter class. The $J^{\text{th}}$ diameter class represents a mature individual that will not grow any larger.
After MPB-induced mortality, trees may die of auxiliary causes. This density-independent mortality is mostly likely to occur from late summer (drought stress) to late winter (ice damage and windthrow). The remaining $(1-m)$ proportion trees grow throughout the spring and summer, automatically entering the next diameter class. Now, the tree density is
$(\text{TPA})_{j+1,y+1}=(1-m)\times(\text{TPA}^{\prime})_{j,y}-(\text{KTPA})_{j,y},\quad j=2,\ldots,J-2.$ (12)
The $J^{\text{th}}$ diameter class receives individuals from the second-largest diameter class, but also retains the mature individuals:
$(\text{TPA})_{J,y+1}=(1-m)\times\left((\text{TPA}^{\prime})_{J-1,y}+(\text{TPA}^{\prime})_{J,y}\right).$ (13)
Trees are recruited into the smallest diameter class, such that there are always $K$ standing trees per acre, which includes live trees and MPB-killed trees that have not yet dropped their needles. We assume that MPB-killed trees do not create light gaps for 3 years, and that all density-independent mortality results in an immediate light gap. With these assumptions, recruitment can be written as
$(\text{TPA})_{1,y+1}=\overbrace{K}^{\text{Carrying capacity}}-\overbrace{\left(\sum_{j=1}^{J}\left(1-m\right)\times(\text{TPA}^{\prime})_{j,y}\right)}^{\begin{subarray}{c}\text{Trees remaining after}\\ \text{density-independent mortality}\end{subarray}}+\overbrace{\left(\sum_{s=y-2}^{y}\sum_{j=1}^{J}(\text{KTPA})_{j,s}\right)}^{\begin{subarray}{c}\text{MPB-killed snags}\\ \text{that still have needles}\end{subarray}}.$ (14)
We assume that each tree’s DBH grows $4\text{mm}=0.157\text{in}$ each year, which is consistent with empirical estimates (Mata,, 2003; Heath and Alfaro,, 1990; McLane et al.,, 2011). However, like all pines, lodgepole pine grows faster in youth and slower in age, so 4mm here is a convenient approximation, representing the average growth increment over a tree’s lifespan.
Although lodgepole pine trees are capable of indeterminate growth, intraspecific competition in dense stands constrains the maximum diameter. The vast majority of lodgepole pine trees in even-aged stands do not exceed 20–24 inches DBH (Rodman et al.,, 2022; Negrón,, 2020; Huang,, 1999), and many U.S. Forest Service reports use a “$>18$ inches” category to describe lodgepole diameters (e.g., Cole,, 1969; Stipe,, 1974; McGregor,, 1973). Therefore, we assume that the trees stop growing after reaching 20 inches DBH. To summarize the fixed-value parameters associated with tree growth, we have $DBH_{1}=0.157$, $DBH_{J}=20-\left(20\bmod 0.157\right)\approx 20$, and $DBH_{j}=DBH_{j-1}+0.157$ for $j=2,\ldots,J$. Assuming that there are no MPB, and that $m$ is large enough that the $J^{th}$ DBH class contains very few individuals, we to can perform recursion on Eq. 12 to approximate the equilibrium distribution of tree diameters as $(\text{TPA})_{j,\infty}=(1-m)^{(j-1)}mK$. This geometric distribution matches the pre-outbreak distribution in the Klein et al., dataset (Fig. 13).
Figure 13: The pre-outbreak diameter distribution of lodgepole pine is effectively modelled by assuming density-independent mortality and instant recruitment (see 12 & Eq. 14). The theoretical distribution of trees per acre in the $j^{th}$ diameter class is $(\text{TPA})_{j,\infty}=(1-m)^{(j-1)}mK$, so the logarithm of trees per acre within should be approximately linear with respect to diameter at breast height (DBH). Data from Klein et al., (1978). The predictions were generated with posterior means of model parameters.
2.4 Model-fitting
The parameters of the simulation model are estimated with three statistical models: the lodgepole-dynamics model, the Allee-effect model, and the outbreak-dynamics model. The combined statistical models are conceptually identical to the simulation model, but their mathematical descriptions are more verbose, and thus they are relegated to Appendix A. All models were fit to data with Stan (Stan Development Team,, 2020), a Bayesian model-fitting software that implements Hamiltonian Monte Carlo.
The first statistical model (Appendix A.2) estimates the demographic parameters of lodgepole pine: density-independent mortality and forest carrying capacity ($m$ and $K$). This model was fit with the 1965 pre-outbreak tree diameter distribution in the Klein et al., dataset (sample size = 7). Our second statistical model (Appendix A.6) describes the Allee effect and was fit with Boone et al., data (sample size = 28). The posterior distribution of the Allee threshold parameter was used to create an informative prior for the Allee effect in our third statistical model.
[…]
In the third statistical model, the Allee threshold parameter is given a half-normal prior, with a standard deviation equal to half of the point estimate of $a$ (the mean of the marginal posterior) from the second model (which estimates the Allee effect in the Boone et al., data). This prior expresses the belief that the Allee threshold in the Boone et al., data is an upper bound on the threshold in the Klein et al., dataset, since the Klein et al., data are averages over a much larger spatial scale (259 hectares vs. 15 hectares) and MPB are known to be highly spatially aggregated. All other parameters are given weakly-informative priors, which are later shown to have negligible influence on the results.
2.5 Causal analysis
When the simulation model is instantiated with realistic parameter values, the five focal life-history traits are turned on. To turn the life-history traits off, we alter specific parameter values.
To turn off the Allee effect, we set the Allee threshold, $a$, to infinity. As a consequence, the probability of a successful attack is always one, or equivalently, the number of initial attacks equals the number of successful attacks. To turn off stand-dependent dispersal, we set $\lambda_{0}$ to infinity, such that the $s_{y}=s_{0}$ and $\lambda_{y}=1$ in Eq. 2 & Eq. 3 respectively. To turn off the size-dependent attack preference, the piecewise linear function for attack preference, Eq. 7, becomes a step function,
$(\text{preference})_{j}=\begin{cases}1&\text{if }\text{{DBH}}_{j}>-\beta_{0}/\beta_{1}\\ 0&\text{otherwise}.\end{cases}$ (15)
All trees above the attack threshold, $-\beta_{0}/\beta_{1}$, are equally attractive. An alternative approach would be to make all trees (susceptible or not) equally attractive, but this approach would vastly increase the number of trees available to the beetles. In doing so, we would be modulating two key ecological factors — size-dependent host-feeding and resource availability — which which diverges further from the real-world scenario and is therefore less causally relevant (Lewis,, 1979)
To turn off size-dependent fecundity, we set $\zeta_{1}=0$ and set $\zeta_{0}$ to the average emergence density (across killed tree surface area in the Klein et al., dataset), which is calculated as
$\sum_{y=1}^{Y^{\prime}}\sum_{j=1}^{J^{\prime}}\begin{cases}\frac{\left(\zeta_{0}+\zeta_{1}\text{DBH}_{j}\right)(\text{KTPA}_{j,y}\times(\text{surface area per tree})_{j})}{\sum_{z=1}^{Y^{\prime}}\sum_{k=1}^{J^{\prime}}(\text{KTPA}_{k,z}\times(\text{surface area per tree})_{k})},&\text{if }\text{DBH}_{j}>-\frac{\zeta_{0}}{\zeta_{1}}\\ 0,&\text{otherwise}.\end{cases}$ (16)
Note here that $J^{\prime}$ and $Y^{\prime}$ are the number of diameter classes and years in the Klein et al., dataset, and should not be confused with the more granular diameter classes and the indeterminate number of years in the simulation model.
Finally, we turn off time-scale separation by forcing lodgepole pine to grow quickly. To accomplish this, we iterate pine mortality, growth, and recruitment (Eq. 12–Eq. 14) 50 times between bouts of beetle attacks, such that year-old trees have 8-inch diameters.
We create simulation scenarios where either 0, 1, or 2 of the life history traits have been turned off. The result is 10 simulation scenarios: 9 counterfactual and 1 realistic (i.e., all of the traits are turned on). For each scenario, we draw 1,000 sets of parameters from the joint posterior distribution of the statistical models. For each parameter set, we simulate 3,000 years of data, starting with zero beetles and equilibrium forest conditions.
2.6 Describing population dynamics
We track 5 key metrics of MPB population dynamics: peak beetle density, length of the outbreak growth phases, outbreak period, large tree density, and proportion of time spent in the endemic state. To compare model predictions to the real world, we also calculated observation-based estimates of these 5 features.
As a preliminary step to quantifying the 5 features, we must distinguish persistent fluctuations (i.e., outbreaks and/or population cycles) from noisy fluctuations. First, we create a smoothed time series by applying a centered moving average with a three-year window. Second, we find peaks and troughs in the smoothed time series; these serve as initial guesses for the true peaks and troughs. Third, we go back to the original time series and identify the true peaks and troughs, e.g., a true trough is the minimum value between successive “guess peaks” identified in the smoothed time series. To prevent sub-critical immigration from being classified as an outbreak, we disregard sets of peaks and troughs where the $B_{\text{peak}}-B_{\text{trough}}<600$; note that $600$ is twice the stochastic immigration parameter $I$ (see Eq. 4).
Two of the metrics require an operational definition of an endemic MPB population. There are multiple definitions in the literature: less than 40 female beetles per ha (Carroll et al., 2006b, ), less than 5 attacked trees per ha (Boone et al.,, 2011), and less than 0.025 infested large-diameter trees per ha (Amman,, 1984). Of these, we select the smallest value ($40$ females/ha $\approx$ $16$ females/acre) to operationalize the endemic state. This choice is based on two considerations: a) smaller values align more closely with aerial survey data, where the majority of stands appear to have no infestation; and b) our simulation model does not explicitly represent endemic beetle populations, leading to a slight underestimate of beetle density.
  1. 1.
    Peak beetle density. For each simulated time series, we take the third quartile of all peak values of $B_{y}$, the density of emerging female beetles. The third quartile represents a typical large outbreak, which is relevant because large outbreaks are likely the driver of range expansion (see the Discussion, Section 4).
    Because beetles or emergence holes are rarely observed across multiple years, the only observation-based estimate of peak beetle density comes from the Klein et al., dataset. This dataset is also used to parameterize the simulation model, so the predictions and observations are not independent.
  2. 2.
    Outbreak growth phase length For each outbreak in a simulated time series times, we calculate the difference between the year of peak beetle density and the year before beetle density first exceeds the endemic threshold of $16$ females/ha. Then, we take the mean of this quantity across all outbreaks.
    An observation-based estimate is calculated using British Columbia’s Aerial Overview Survey data from 1990-2021. First, we estimate the number of infested trees using the methodology in Fig. D.3 caption, aggregate data into $1\text{km}^{2}$ pixels, and subset to locations with moderate-to-severe outbreaks (cumulative infestations in the upper quartile). Next, we calculate the difference between the year with the maximum number of infested trees and the year before infested trees were first detected. Finally, we take the mean of this quantity across all locations.
  3. 3.
    Outbreak period. This quantity is simply the mean number of years between peaks (whether cycles or outbreaks) in the simulated time series. The observation-based estimate was sourced from studies that have reconstructed region-wide outbreaks with tree ring data (Alfaro et al.,, 2003; Negrón and Huckaby,, 2020; Jarvis and Kulakowski,, 2015; Alfaro et al.,, 2009; Axelson et al.,, 2009). More precisely, the observation-based estimate is the median (across studies) of the mean outbreak recurrence time.
  4. 4.
    Large tree density. The large tree density is simply the temporal average of trees per acre with $DBH\geq 8$ inches. Here, there are two observation-based estimates. The first comes from the Klein et al., dataset, and is therefore not independent of the simulation-based predictions. The second is the median (across several studies) of trees per acre with $DBH\geq 8$ inches, prior to a MPB outbreak (McGregor,, 1973; Gibson and Bennett,, 1978; Hamel et al.,, 1975; Stipe,, 1974, 1976; Mitchell and Preisler,, 1991; Amman,, 1969; Negrón,, 2020; Evenden and Gibson,, 1940). This is an overestimate of the temporal average of large-tree density, but it may be a sensible approximation, given that MPB populations are usually in the endemic state.
  5. 5.
    proportion of time spent in the endemic state. This is straightforwardly computed as the proportion of years with less than 16 emerging females per acre. The observation-based estimate comes from British Columbia’s Aerial Overview Survey data (years: 1962–2021). Specifically, we aggregated data into $1\text{km}^{2}$ pixels, subsetting to locations with moderate-to-severe outbreaks (cumulative infestations in the upper quartile), and computed the proportion of years where no infestations were detected.
The five metrics above can be used heuristically to classify simulated time series into qualitative classes of population dynamics: episodic boom-bust dynamics, population cycles, or equilibrium dynamics. If the proportion of time spent in the endemic states is $\geq 1/2$, then the time series has Boom-bust dynamics. If the proportion of time spent in the endemic states is $<1/2$, and the outbreak period is $<50$ years, then the time series has Cycles. If the proportion of time spent in the endemic states is $<1/2$, and the outbreak period is $\geq 50$ years, then the time series has Equilibrium dynamics.
3 Results
The statistical model successfully predicts in-sample data (Fig. 14, panel A) and out-of-sample data. The exception is that the out-of-sample predictive intervals for 1970 are wide (Fig. 14, panel B), which upon further investigation, was found to be the result of uncertain estimates of the stand-dependent dispersal parameters, $\lambda_{0}$ and $\lambda_{1}$ (Fig. 14, panel C). The estimation of these parameters is highly dependent on the data from 1969–1970. This would be indicative of overfitting (i.e., fitting the noise) if the decrease in attack density in 1970 was the result of noise, rather than stand-dependent dispersal. The obvious source of statistical noise is the weather, but the weather leading up to the summer of 1970 was unexceptional (Figs. D.5 & D.4). Taken together, evidence for the presence of stand-dependent dispersal (Section 1.2.2) and the absence of anomalous weather patterns imply that the model did not overfit.
Various diagnostics showed that the statistical models converged to the global posterior distribution, and that nearly all prior distributions had a negligible impact on the results (Appendix B.1). The only exception was the Allee threshold parameter, whose marginal posterior variance was similar in magnitude to its marginal prior variance, indicating that the Klein et al., data is not informative with respect to the Allee effect. This is unsurprising, as discerning the Allee effect is challenging without data from trees that were attacked but not successfully killed (see Section Section 1.2.1), data which is absent in Klein et al.,’s 1978 publication. However, the Allee threshold estimate is not devoid of information, since the prior distribution was derived from the Allee effect sub-model fit to the Boone et al., dataset (see Table C.2 & Appendix A.6).
Figure 14: Observations of successful attack density in the Klein et al., dataset (operationalized as gallery starts), with superimposed model-based predictive intervals (PIs). Panel A shows in-sample predictions for the full statistical model; panel B shows out-of-sample, leave-one-year-out (LOYO) predictions, where the model is not trained on the focal year’s data; and panel C shows LOYO predictions when the stand-dependent dispersal parameters ($\lambda_{0}$ and $\lambda_{1}$) are fixed at the medians of their respective marginal posterior distributions (see Table 1). Panels B & C demonstrate that data from a single beetle generation (1969–1970) has an overwhelming influence on the estimation of $\lambda_{0}$ and $\lambda_{1}$, but that all other parameter estimates are robust.
The statistical models also produced reasonable parameter estimates. In particular, the posterior distributions of attack density and emergence density (from a 10-inch diameter tree) are consistent with previous estimates (Figures 15 & 16). The mean overstory density, 338 trees per acre or 835 trees per hectare, agrees with estimates from regenerating stands in southern British Columbia (Nigh et al.,, 2008, mean $=$ 893 stems/ha) and lodgepole-dominated stands in the U.S. Rocky Mountains (Pfeifer et al.,, 2011; Meddens et al.,, 2011, 2012; mean $=$ 1065 stems/ha). The minimum susceptible tree diameter is estimated at 7.34 inches, which is consistent with the often-cited 8-inch (or 20 cm) threshold (Hopping and Beall,, 1948; Cole,, 1969; Safranyik,, 1988). The baseline dispersal mortality (i.e., 1-$s_{0}$) is 0.55, which is consistent with the 0.5–0.7 range provided by previous studies of bark beetles (Schmid,, 1977; Safranyik,, 1989; Latty,, 2007; Pope et al.,, 1980).
Figure 15: Frequency distributions of attack density. The Model posterior category gives the marginal posterior of $1/\gamma$ from our statistical model of the Klein et al., outbreak. The AB disk data category contains the within-site means of attack density from 300 different sites across Alberta (author N/A,, 2020). The Other studies category includes estimates obtained via entry holes or gallery starts from a number of studies, mostly United States Forest Service report: Tishmack et al., (2005); Mccambridge, (1967); Negrón, (2018); Reid, (1963); Raffa and Berryman, (1983); Rasmussen, (1980); Shepard, (1965); Knight, (1959); De Leon, (1939); Safranyik, (1988); Safranyik and Vithayasai, (1971); Safranyik, (1968); Parker, (1979); Schmid, (1972); Blackman, (1931); Whiteside, (1937); Washburn and Cole, (1959); author N/A, (1963); Peterman, (1974); Klein et al., (1978); Safranyik and Linton, (1991).
Figure 16: Frequency distributions of female emergence density. Females are estimated with (2/3)×(3/2)×(2/3$\times$(3/2$\times$( 2 / 3 ) × ( 3 / 2 ) × emergence holes, or with (2/3)×(2/3$\times$( 2 / 3 ) × emerging beetles or spring brood size. Note that (2/3) is the typical proportion of females in a brood (Reid,, 1962; Cole et al.,, 1976), and (3/2) is the average number of beetles emerging from each hole (Safranyik and Linton,, 1985; Peterman,, 1974, Fig. 13). The Model posterior category gives the marginal posterior of $\zeta_{0}+\zeta_{1}*10$, the emergence density on a 10-inch DBH tree; 10 inches is approximately the average DBH of MPB-killed trees in the Klein et al., dataset. The AB disk data category contains estimates of emergence from 300 different sites across Alberta (author N/A,, 2020). The Other studies category includes estimates from a number of studies, mostly United States Forest Service reports: Tishmack et al., (2005); Negrón, (2018); Reid, (1963); Raffa and Berryman, (1983); Rasmussen, (1980); De Leon, (1939); Safranyik, (1988); Cole, (1969); Schmid, (1972); Whiteside, (1937); Beal, (1938); Peterman, (1974); Klein et al., (1978); Safranyik and Linton, (1985, 1991).
Parameter Short description Mean SD $\text{CI}_{2.5\%}$ $\text{CI}_{97.5\%}$
a Allee threshold 392 331 12.1 1211.68175
$s_{0}$ max dispersal survival 0.474 0.0676 0.363 0.627
$\lambda_{0}$ dispersal threshold 0.652 0.0339 0.582 0.705
$\lambda_{1}$ dispersal sensitivity 49.7 96.6 17.2 262
[…]
K forest carrying capacity 904 313 413 1630.37825
m annual tree mortality 0.0518 0.00637 0.0376 0.0634
Derived quantities Short description Mean SD $\text{CI}_{2.5\%}$ $\text{CI}_{97.5\%}$
$-\beta_{0}/\beta_{1}$ susc. DBH threshold 7.43 0.643 6.04 8.43
$1/\gamma$ attack density 66.2 3.78 59.2 74.1
$\zeta_{0}+\zeta_{1}*10$ emerge density (10 in DBH) 244 25.4 192 297
$K-\sum_{j=1}^{18}(1-m)^{(j-1)}mK$ Overstory density 338 86.0 193 529
Fixed parameter Short description value
$P_{I}$ Prob. immigration 0.05
$I$ immigrating females 300
$\sigma_{\text{env}}$ Environmental stochasticity 0.2
Symbol N/A max DBH 20
Symbol N/A annual DBH growth 0.157
Table 1: Summary statistics of marginal posterior distributions for model parameters, including the Mean, standard deviation, and credible intervals ($\text{CI}_{2.5\%}$, $\text{CI}_{97.5\%}$). The table also includes derived quantities, including the diameter at breast height (DBH) of the smallest susceptible trees; the model-based estimate of successful attack density (gallery starts per square meter); the emergence density (females per square meter) for 10-inch DBH tree; and the density of overstory trees (stems per acre). Overstory trees are operationalized with a standard cutoff: DBH $\geq 7\text{cm}\approx 2.76\text{in}$. This explains the summation from 0 to 18 in the formula: 2.76in cutoff0.157in per $\frac{2.76\text{in cutoff}}{0.157\text{in per year}}\approx 18\;\text{years}$. Finally, the table contains fixed parameters, whose values were obtained via literature search. For more elaborate descriptions and parameter units, see supplementary Table A.1. For the summary statistics of sub-model noise parameters, see supplementary Table C.1.
Given our understanding of MPB dynamics, three life-history traits are necessary for episodic outbreaks: Stand-dependent dispersal, an Allee effect, and time-scale separation (Fig. 19; Table 2, column Boom-Bust). Without the stand-dependent dispersal, MPB and lodgepole pine attain a noisy equilibrium where there is a consistently low density of beetles (Fig. 19C), but not low enough to technically be endemic. Without the Allee effect, a classic population cycle appears (Fig. 19D). Without time-scale separation, MPB and lodgepole pine attain a noisy equilibrium; in contrast with the no-stand-dependent-dispersal scenario, this equilibrium sustains large numbers of beetles and susceptible trees (Fig. 19E).
The inclusion of stand-dependent dispersal leads to better predictions of outbreak magnitude and large-tree density (Fig. 17). The Allee effect leads to better predictions of the outbreak recurrence time and growth phase length. Specifically, without the Allee effect, the cycles occur too frequently and the growth phase of each cycle is too long. Stand-dependent dispersal primarily explains the fact that large trees remain despite MPB outbreaks.
The two remaining life-history traits — size-dependent tree preference and size-dependent fecundity — have no effect on the qualitative nature of population dynamics, i.e., whether a simulation was classified as Boom-Bust or Cycles. However, they do have quantitative effects on some of the fine-grained properties of outbreaks: the combination of both size-dependent tree preference and size-dependent fecundity lead to a 25% increase in the peak beetle density (from 16 to 20 thousand) and a 10% decrease in the length of growth phase (from 5.8 years to 5.2 years). In other words, the combination of these life-history traits makes outbreaks moderately more explosive and slightly more rapid.
    Life history traits     Has stand-dependent dispersal     Has Allee effect     Has size-dependent preference     Has size-dependent fecundity     Has time-scale separation     Metrics Mean $\pm$ (SD)     Peak beetle density, 1000s of emerging females per acre     Outbreak growth phase, years     Outbreak period, years     Large tree density, trees $\geq$ 8 in DBH per acre     Prop. time endemic     Prop. sims without peaks     Prop. sims without outbreaks     Probability of Dynamics     Boom-Bust     Cycles     Equilibrium
20 (8.7) 5.2 (1.1) 44 (130) 44 (9.6) 0.66 (0.16) 0.00 (0.00) 0.00 0.86 0.10 0.03
9.5 (10) 7.2 (1.1) 179 (251) 10 (12) 0.16 (0.24) 0.01 (0.09) 0.44 0.16 0.05 0.79
6.6 (3.2) 9.4 (2.3) 28 (63) 38 (9.9) 0.07 (0.15) 0.00 (0.00) 0.04 0.05 0.83 0.13
16 (7.2) 5.8 (1.4) 51 (57) 43 (9.3) 0.66 (0.16) 0.00 (0.00) 0.00 0.87 0.09 0.03
16 (6.7) 5.8 (1.3) 60 (184) 45 (10) 0.65 (0.16) 0.01 (0.08) 0.01 0.86 0.10 0.03
73 (24) NA 4.8 (0.57) 46 (11) 0.00 (0.03) 0.00 (0.00) 0.99 0.00 1.0 0.00
3.3 (0.88) 15 (1.8) 278 (378) 2.4 (4.2) 0.00 (0.00) 0.01 (0.09) 1.0 0.00 0.06 0.94
11 (13) 8.5 (1.3) 151 (168) 11 (14) 0.18 (0.26) 0.00 (0.06) 0.43 0.18 0.07 0.75
4.5 (2.2) 9.2 (2.2) 1078 (838) 3.8 (8.9) 0.03 (0.13) 0.08 (0.27) 0.87 0.03 0.00 0.97
71 (17) NA 4.7 (0.12) 44 (8.6) 0.00 (0.00) 0.00 (0.00) 1.0 0.00 1.0 0.00
5.6 (1.9) 10 (2.6) 34 (124) 37 (9.5) 0.06 (0.12) 0.00 (0.00) 0.04 0.02 0.78 0.21
5.9 (2.6) 10 (2.7) 34 (127) 39 (11) 0.05 (0.13) 0.00 (0.00) 0.16 0.03 0.78 0.19
77 (17) NA 4.7 (0.14) 47 (8.4) 0.00 (0.00) 0.00 (0.00) 1.0 0.00 1.0 0.00
15 (6.8) 5.8 (1.3) 57 (110) 43 (9.8) 0.66 (0.17) 0.00 (0.00) 0.00 0.87 0.08 0.05
71 (13) NA 4.2 (0.33) 41 (6.8) 0.01 (0.09) 0.01 (0.10) 1.0 0.01 0.99 0.00
68 (16) NA 4.9 (0.19) 46 (9.2) 0.01 (0.09) 0.01 (0.10) 1.0 0.01 0.99 0.00
Table 2: Results of the simulation study. A check mark indicates that a life history trait is present in the simulation scenario. The quantitative values take the format of “Mean $\pm$ (SD)”, where both summary statistics are calculated across the posterior distribution. In addition to the 5 metric of population dynamics (presented in Section 2.6), there are two extra metrics: 1) the proportion of simulations where neither peaks nor troughs were detected after applying the moving-average smoothing and various thresholds described in Section 2.6; 2) the proportion of simulations where no outbreaks were detected, where an outbreak is defined as a trough $\rightarrow$ peak $\rightarrow$ trough sequence where the population attains the endemic state ($<16$ females per acre) before and after the peak. If a simulation scenario produced both NA and non-NA values for a particular metric, then the NA values were omitted when calculating that metric.
Figure 17: A comparison of population dynamic metrics for various combinations of stand-dependent dispersal and the Allee effect. The histograms show the frequency of predicted values across posterior samples. The vertical lines show observation-based estimates from the Klein et al., dataset or independent sources. Simulation scenarios without stand-dependent dispersal resulted in few peaks and outbreaks (see Table 2), hence the “N/A” label. The five metrics and the observation-based estimates are explained in Section 2.6.
Figure 18: The posterior probability of various types of population dynamics, across various simulation scenarios where stand-dependent dispersal and the Allee effect have been turned on or off.
Figure 19: A time series of Aerial Overview Survey data, along with representative time series from various simulation scenarios.
4 Discussion
Stand-dependent dispersal, an Allee effect, and time-scale separation are jointly necessary for realistic mountain pine beetle outbreaks. Time-scale separation (i.e., trees grow slower than beetles) ensures that killed trees cannot be immediately replaced by new mature pines, thus preventing equilibrium dynamics. Stand-dependent dispersal (i.e., beetles disperse farther through sparse forests) leads to high levels of MPB mortality, and the Allee effect finishes the MPB population off, bringing it to the endemic state. These two inhibitory processes (Stand-dependent dispersal and the Allee effect) explain why MPB populations spend the vast majority of time in their endemic state, and why subsequent outbreaks are so explosive: with MPB populations suppressed, large trees accumulate, providing proverbial fuel for the beetle fire.
The process of stand-dependent dispersal explains why beetles don’t overexploit trees. It is perhaps trivial to say that if beetles disperse away and die before they have killed all the trees, then the beetles won’t kill all the trees. However, the lack of over-exploitation, or equivalently, the seemingly premature “bust” phase of the “boom-bust cycle”, is a ubiquitous feature of outbreaks (Fig. 5) and is thus worth connecting to MPB life-history. This also suggests an optimistic perspective: outbreaks are much less destructive than they theoretically could be. Stand-dependent dispersal probably occurs because chemo-sensory constraints prevent beetles from aggregating on small trees (Section 1.2.2), even though beetles have positive fitness when they do successfully attack small trees. Without stand-dependent dispersal, the system would reach a stochastic equilibrium, creating a forest with few large pine trees.
The Allee effect explains why outbreaks are so extreme when they do occur. The Allee effect suppresses MPB populations, keeping them in their endemic state until enough large trees have accumulated that a small inoculum of beetles — either from long-distance dispersal or a stochastic successful attack –– can grow explosively. Without the Allee effect, lodgepole pine and MPB enter a predator-prey cycle.
Classic predator-prey theory predicts that long-lived prey should make predator-prey cycles unlikely. One might naively expect that the time-scale separation between beetles and trees would allow MPB to grow rapidly enough to prevent an unchecked proliferation in pine abundance, thus preventing the prey-escape phase of a population cycle (De Roos et al.,, 1990) wherein the prey species can grow unimpeded by the predator. Even though pines live much longer than beetles, we still see population cycles in the counterfactual simulations (when the Allee effect is turned off). This can be explained by the phenomenon of cohort resonance (Bjørnstad et al.,, 2004). During an outbreak, many pines are killed at roughly the same time. Subsequently, pines are recruited at the same time, and then become susceptible to MPB at the same time. Thus, there are a brief few years where the pines that MPB can “see” are growing faster than exponentially.
Our model accurately predicts the return time of MPB outbreaks, which is somewhere between 20–50 years according to tree-ring-based reconstructions. Our counterfactual simulations with no Allee effect, resulting in classic population cycles, predict an outbreak period that is too short (approximately 20 years). Other models in the literature claim an outbreak periodicity that is too long (Powell and Bentz,, 2009; Duncan et al.,, 2015; Brush and Lewis,, 2023).
The period of MPB outbreaks is primarily determined by the time it takes lodgepole pine to grow into a susceptible age class. Thus, if MPB were to kill all of the trees, then outbreaks would occur approximately every 55 years: 50 years for the trees to become susceptible and several more years for the outbreak and subsequent recruitment of pine saplings. However, there are two real-world complications. First, stand-dependent dispersal results in a large number of medium-sized trees surviving the outbreak, so the observed periodicity of 35 years partially reflects the time it takes for medium-sized trees to become large trees, e.g., it takes approximately 25 years for an 8-inch tree to become a 12-inch tree. The 2000s outbreak in British Columbia killed more medium-sized trees than usual, suggesting that it will take longer than usual for the next outbreak to occur. Second, large trees are a pre-requisite for outbreaks, but widespread outbreaks require widespread perturbations that allow several populations to synchronously break out of the low-density endemic state. This perturbation is usually realized as drought conditions. Tree growth and climate patterns lead to a right-skewed distribution of waiting times between outbreaks, where the minimum time is determined by tree growth, and the shape of the long right tail is determined by drought probability.
The life-history processes of size-dependent fecundity and size-dependent attack preference do not explain the qualitative behavior of MPB dynamics, but they do explain about 25% of the magnitude of outbreaks (i.e., the peak number of beetles). Size-dependent attack preference focuses the beetles on the large trees. Under size-dependent fecundity, the large trees are the most productive (from the beetles’ perspective), so more beetles are produced before most of the susceptible trees are eaten up, triggering stand-dependent dispersal and the subsequent “bust” phase of the outbreak.
The peak number of beetles is a highly relevant statistic, since it is likely the main driver of MPB range expansion. Only a minuscule fraction of emerging beetles undergo long-distance dispersal (estimated at 0–2%), so outbreak expansion via long-distance dispersal requires severe local infestations, such that emigrating beetles are numerous enough to mass attack trees — and in doing so, overcome the Allee effect — at their destination. In other words, range expansion requires that peak beetle density exceeds the donor threshold (terminology introduced by Johnson et al.,, 2006). The concept of a donor threshold may explain why outbreaks are spatially correlated at greater distances in British Columbia than in the western United States (U.S. pine forests generally have less pine biomass $\rightarrow$ less peak beetles), and why MPB spread into central Alberta via long-distance dispersal from British Columbia, rather than slowly spreading up the east side of the Rocky Mountains.
Through our model-building process and our simulation study, we have gathered insights about the determinants of MPB’s distinctive population dynamics. Stand-dependent dispersal and the Allee effect are essential, whereas size-dependent attack preference and size-dependent fecundity are less important. Beetle attack density may increase throughout an outbreak, but this has a negligible effect on population dynamics (Appendix B.2). MPB’s preference for large-diameter trees may become exacerbated when there are many beetles, but this has a negligible effect on population dynamics, at least in the epidemic phase of an outbreak (Appendix B.1).
Our research offers a revised perspective on the density-dependent dynamics of MPB, challenging established assumptions. In previous work, the number of new infestations is assumed to be proportional to both beetles and trees (e.g. Burnell,, 1977; Heavilin et al.,, 2007; Křivan et al.,, 2016), as in the classic Lotka-Volterra model. Yet, our analysis reveals a different perspective for cases of intermediate-to-high tree density (i.e., before stand-dependent dispersal kicks in): the number of new infestations is only proportional to beetles, reflecting the fact that outbreaking beetles are limited by their ability to overcome tree defenses, not by their ability to find susceptible trees. While overcompensatory density dependence has been a common assumption in past studies (e.g. Goodsman et al.,, 2017, 2018; Lewis et al.,, 2010; Duncan et al.,, 2015), we argue that MPB experiences near-perfect compensatory density dependence (see Section 1.2.2). Previous research has assumed that the probability of a successful mass attack depends on the quotient of beetle and tree density (e.g., Burnell,, 1977; Goodsman et al.,, 2016; Brush and Lewis,, 2023). Instead, we propose that beetle density alone is a more significant determinant of the success probability, as detailed in Appendix B.3.
Putting all of these results together, we can simplify our simulation model (Section 2.3) to formulate a minimal mechanistic model of MPB dynamics. Because size-dependent preference and size-dependent fecundity are non-essential, we drop explicit age structure and only track the density of all susceptible trees (i.e., trees with DBH $>$ 8 inches). All susceptible trees are equally preferred for mass attacks, and all killed trees give rise to the same number of beetles. For simplicity, stochasticity is eliminated. The minimal mechanistic model can be written as
$\overbrace{A_{y+1}}^{\mathclap{\text{Attacking females per acre}}}=A_{y}\times\underbrace{(1-e^{\frac{-2.3A_{y}}{a}})}_{\mathclap{\text{Allee effect}}}\times\overbrace{\left(\zeta^{*}\;\gamma^{*}\right)}^{\mathclap{\raisebox{6.02777pt}{\scriptsize\text{Emerging females per succ. attack}}}}\times\underbrace{s_{0}\left(1+\exp\left(-\lambda_{1}\left(\frac{S_{y}^{\prime}}{S_{max}}-\lambda_{0}\right)\right)\right)^{-1}}_{\mathclap{\text{Survival after stand-dependent dispersal}}}+\overbrace{\epsilon}^{\mathclap{\raisebox{6.02777pt}{\scriptsize\text{Endemic beetles}}}}$ (17)
$\underbrace{S_{t+1}}_{\mathclap{\text{Susceptible trees per acre}}}=\overbrace{g(K^{*}-S_{y})}^{\mathclap{\text{Recruitment of susceptible trees}}}-\underbrace{A_{y}\times(1-e^{\frac{-2.3A_{y}}{a}})\times\overbrace{\gamma^{*}}^{\mathclap{\text{Trees killed per succ. beetle attack}}}}_{\mathclap{\text{Trees killed by MPB}}}.$ (18)
Here, $A_{y}$ is the female attack density in the $y^{\text{th}}$ year, $S_{y}$ is the density of susceptible trees, $S_{y}^{\prime}$ is the density of susceptible trees remaining after beetle attack (i.e., $S_{y}^{\prime}=S_{y}-A_{y}\times(1-e^{\frac{-2.3A_{y}}{a}})$), the parameter $\zeta^{*}$ is the average emerging females per killed tree, $\gamma^{*}$ is the average number of killed trees per successful attack, $K^{*}$ is redefined as the carrying capacity of the susceptible trees, $\epsilon$ is the now-explicit number of endemic female beetles, $g$ modulates the rate of recruitment into the class of susceptible trees, and all other parameters have their original meanings (see Table A.1).
The minimal model has several purposes. 1) It is descriptive, compactly representing the density-dependent dynamics of MPB and lodgepole pine. 2) It demonstrates that there are two stable equilibria: an endemic equilibrium with few beetles, and an epidemic equilibrium with many beetles. This bistability is typical of systems with a strong Allee effect (Taylor,, 2012), so it is no surprise that the endemic equilibrium disappears when the Allee effect is turned off (Fig. 20C) 3) It shows the MPB/lodgepole system is an instance of an excitable dynamical system, where perturbations can lead to long excursions in phase space, before returning to a resting state (Fig. 20A). Typically, excitable systems have a single stable equilibrium (Strogatz,, 1994, pg. 118), whereas our system has two. However, the endemic equilibrium is globally-attracting as a practical matter, since the epidemic equilibrium has a minuscule basin of attraction (Fig. D.6).
Figure 20: Phase portraits of the minimal mechanistic model (Eq. 17 & Eq. 18), for scenarios where stand-dependent dispersal and the Allee effect have been turned on or off. Most model parameters were set to their respective posterior means (see Table 2). The exceptions are $\gamma^{*}=1/360$ and $\zeta^{*}=1300$, which are approximately the respective means (across trees) of attacks and emergences in the Klein et al., dataset. The carrying capacity’s value, $K^{*}=130$, is approximately the density of trees with $\text{DBH}\geq 8\text{in}$ in the Klein et al., dataset. The recruitment parameter’s value, $g=0.036$, was obtained by solving the differential equation $dS(t)/dt=g(K-S(t))$ with initial condition $S(0)=0$, then solving for the $g$ where 90% of carrying capacity is attained in $t=50$ years; lodgepole pine grows attains an 8-inch DBH and becomes susceptible after approximately 50 years.
While our models thoroughly describe the density-dependent dynamics of MPB, they do not account for the impacts of environmental heterogeneity or stratified dispersal. There are additional un-modelled processes, including temperature-dependent phenology, beetle vigor depending on natal tree size, competition with other bark beetle species, and predation by natural enemies like woodpeckers and parasitoid wasps. While we are confident in the structure of our model, parameter estimates primarily rely on data from a single outbreak at a single geographic location. Thus, our models should not be used to make precise quantitative predictions.
Through the development and analysis of our age-structured simulation model, we have shed light on the life-history traits underlying outbreaks, and challenged established ideas in the MPB modelling literature. Our findings are succinctly summarized in a minimal mechanistic model, which incidentally establishes the MPB/lodgepole system as one of the first empirical examples of an excitable dynamical system in ecology. This work underscores the importance of detailed case studies and looking beyond traditional predator-prey models to understand complex population dynamics. Future research should leverage our description of dynamics to model the extensive Aerial Overview Survey data from the United States and British Columbia, both to forecast outbreaks and to understand large-scale dispersal patterns.
5 Acknowledgements
We would like to thank Micah Brush and Antonia Musso for insightful discussions; and Xiaoqi Xie for obtaining the Biosim data (Figs. D.5 & D.4. This research was funded through a grant to the TRIA-FoR Project to ML from Genome Canada (Project No. 18202).
6 Data and code availability statement
All code and almost all data will be made available shortly, both on Github and a TBD data repository.
When seeking data from publications, we initially searched for downloadable data files. Where these weren’t available, we used the GraphGrabber software and Microsoft Excel’s digitizing features to extract data from figures and tables. Fixed-wing Aerial Detection Survey data for the western United States was downloaded from various (region-specific) USDA Forest Service websites. Fixed-wing Aerial Overview survey data for British Columbia was downloaded from the BC forestry FTP site (as of December 12, 2023, https://www.for.gov.bc.ca/ftp/). Heli-GPS Aerial Survey data for Alberta — here utilized in Figs. 1 & 2 — was provided by the provincial government of Alberta. However, this data cannot be made publicly available per our data-sharing agreement.
Appendix A Statistical model description
A.1 Overview
There are three statistical models. 1) The tree growth model (Appendix A.2); 2) The simple Allee effect model (Appendix A.6), which is fit to the Boone et al., dataset; and the complex outbreak model (Appendices A.3A.7), which is fit to the Klein et al., dataset. Table A.1 describes all relevant notation, including subscripts, data, intermediate quantities, and bonafide parameters.
In the complex outbreak model, each life-history process is described in its own sub-model (Appendices A.3A.6). Going further, the deterministic skeletons of the sub-models are strung together to form the integrator model (Appendix A.7). This model integrates over unobserved quantities (e.g., beetles remaining after dispersal) to estimate additional model parameters.
In the main text, we presented the simulation model with each time-step beginning and ending with the density of emerging beetles. This approach was chosen for two reasons: Firstly, flying beetles are a noticeable aspect of actual outbreaks. Secondly, emergence ties directly into our discussion of the interaction between peak beetle density and long-distance dispersal.
In the outbreak statistical model, each time step begins and ends with beetles that have successfully mass attacked a tree. This model structure gives us an additional year of data at the beginning of the outbreak; the later years of the outbreak are dominated by stand-dependent dispersal, and are therefore less informative with respect to the estimation of many non-dispersal parameters.
A.2 Lodgeple pine forest dynamics
We begin by deriving tree ages in a stable population, undisturbed by mountain pine beetle. Let $K$ be the carrying capacity (units: trees/acre). Let $m$ be the annual probability of density-independent mortality, constant across all ages. When a tree dies, a seedling is immediately recruited in its place, such that there are always $K$ trees per acre. The asymptotic density of trees of age $t$ is given by:
$N_{t}=Km(1-m)^{t}$ (19)
Here, $K\times m$ is the density of recruits into the first age class, and $(1-m)^{t}$ is the probability that a tree survives up to age $t$.
The Klein et al., dataset does not contain tree ages, but it does contain densities of trees within diameter classes. To convert diameter to age, we divide DBH (in inches) by 0.15748, the average annual diameter increment of lodgepole pine.
[…]
$\mu_{j,\text{tree}}=\sum_{t=\text{(lower age)}_{j}+1}^{\text{(upper age)}_{j}}Km(1-m)^{t}$ (20)
Finally, the likelihood for the observed tree density, denoted, $(\text{TPA})_{j}$, is given by the normal PDF:
$\log\left(\text{TPA}\right)_{j}\sim\text{Normal}(\log(\mu_{j,\text{tree}}),\sigma_{\text{tree}}).$ (21)
Residual errors in the densities of small and large-diameter trees are made commensurate through the logarithmic transformation. The noise parameter is $\sigma_{\text{tree}}$. The model provides a good fit to the data (Fig. A.1), though there is substantial uncertainty due to the high quotient of parameters to data (3 and 7 respectively).
Figure A.1: Observations vs. predictions for the lodgepole pine forest model. The error bars give the 95% predictive interval, representing both parameter and sampling uncertainty. The dashed grey line represents perfect agreement between observed and predicted values. Data from Klein et al., (1978).
A.3 Size-dependent preference sub-model
We define mass attack preference for trees in the the $j^{th}$ diameter class, in the $y^{th}$ year, as
$(\text{{preference}})_{j,y}=\frac{(\text{{proportion of successful attacks received}})_{j,y}}{(\text{{proportion of total tree surface area}})_{j,y}}.$ (22)
High values of preference indicate that beetles like to attack a DBH class disproportionately, in relation to the surface area that is “owned” by that DBH class. The preference variable can be calculated directly from the Klein et al., data. The data show that beetle preference increases linearly with tree diameter, and that there is clear heteroskedasticity that is explained by small sample sizes for larger DBH classes (Fig. 10)
Preference is normally distributed with mean $\mu_{j,\text{{pref}}}$ and standard deviation $\sigma_{j,y,\text{{pref}}}$:
$(\text{{preference}})_{j,y}\sim\text{Normal}(\mu_{j,\text{{pref}}},\sigma_{j,y,\text{{pref}}}).$ (23)
The mean is a piecewise linear function,
$\mu_{j,\text{{pref}}}=\begin{cases}\beta_{0}+\beta_{1}\cdot\text{{DBH}}_{j}&\text{if }\text{{DBH}}_{j}>-\beta_{0}/\beta_{1}\\ 0&\text{otherwise}.\end{cases}$ (24)
Above, we assume that $DBH_{j}$ is the diameter at breast height for the middle of a diameter class, e.g., if we are considering all trees where $DBH\in[10,12)$, then $DBH_{i}$ in the above equation is 11 inches. In the Klein et al., dataset, there are six 2-inch DBH classes. In the simulation model, each tree age is associated with a unique DBH, separated by intervals of 0.157 inches.
The standard deviation of preference is
$\sigma_{j,y,\text{{pref}}}=\sqrt{\frac{{\sigma_{\text{{pref0}}}^{2}}}{(\text{total surface area})_{j,y}}+\sigma_{\text{{pref1}}}^{2}},$ (25)
where $\sigma_{\text{{pref0}}}$ accounts for a baseline level of variability, and $\sigma_{\text{{pref1}}}$ accounts for increasing variability with smaller sample sizes, with the typical inverse square root scaling that would be expected under the central limit theorem. We do not know the exact sample sizes, so we treat the total tree surface area (within the $j^{th}$ diameter class) as a proxy for sample size.
A.4 Converting beetle attacks to tree deaths
We model killed trees as a linear regression through the origin, with a dynamic noise term to account for some obvious heteroskedasticity (Fig. 12). The killed trees per acre, denoted $(KTPA)_{j,y}$ is normally distributed:
$(\text{{dead}})_{j,y}\sim\text{Normal}(\mu_{j,y,\text{{dead}}},\sigma_{j,y,\text{{dead}}}).$ (26)
The mean and standard deviations are given by the equations,
$\mu_{j,y,\text{{dead}}}=\gamma\frac{(\text{succ. attacks per acre})_{j,y}}{(\text{surface area per tree})_{j}},\quad\text{and}$ (27)
$\sigma_{j,y,\text{{dead}}}=\sigma_{\text{{dead0}}}\frac{(\text{succ. attacks per acre})_{j,y}}{(\text{surface area per tree})_{j}},$ (28)
We fit variants of the above model where the standard deviation increased as a square of the independent variable, where the standard deviation was constant, and where the standard deviation contained both constant and dynamics parts (as in the Eq. 25). The model presented above was the best according to Leave-One-Out Cross Validation with expected log predictive density.
A.5 Size-dependent fecundity
Figure A.2: Observations vs. predictions for the size-dependent fecundity sub-model. The error bars give the 95% predictive intervals, representing both parameter and sampling uncertainty. The dashed grey line diagonal line represents perfect agreement between observed and predicted values. Data from Klein et al., (1978).
The Klein et al., dataset does not contain DBH-specific emergence data. The total number of emergence holes per acre is normally distributed
$(\text{{exit density}})_{y+1}\sim\text{Normal}(\mu_{y,\text{{exit}}},\sigma_{\text{{exit}}}),$ (29)
where the mean is a sum over DBH classes:
$\mu_{y+1,\text{exit}}=\sum_{j=1}^{J}\begin{cases}\overbrace{\left(\zeta_{0}+\zeta_{1}\cdot\text{DBH}_{j}\right)}^{\text{Emergence density}}\cdot\overbrace{(\text{KTPA})_{j,y}\cdot(\text{surface area per tree})_{j}}^{\text{surface area of killed trees}},&\text{if }\text{DBH}_{j}>-\frac{\zeta_{0}}{\zeta_{1}}\\ 0,&\text{otherwise}.\end{cases}$ (30)
Emergence occurs after the MPB brood overwinters, so the calendar year index for the response variable and the mean prediction is $y+1$. The standard deviation is simply a constant. Refining the error structure would require unjustifiable assumptions, given the absence of DBH-specific emergence data. The model predictions are in agreement with the data (Fig. A.2).
A.6 Allee effect
The Boone et al., dataset contains two pertinent variables: density of trees attacked and proportion of attacked trees that are killed. As a first step, we convert the number of attacked trees to the number of attacking beetles. This is done primarily for practical reasons, as we aim to model the Klein et al., data, which lacks information on the number of attacked trees. Additionally, from a biological standpoint, using the number of attacked trees as a success indicator could be misleading. For instance, a scenario where numerous beetles attack a single large tree would erroneously imply a low success rate. This contradicts observations that beetles are more successful when they concentrate on larger trees.
To estimate the number of attacking beetles, we multiply the attacked trees by 360, the average number of attacks per successfully attacked tree in the Klein et al.,. Since successfully attacked trees receive more attacks, this is probably an overestimate of beetles per attacked tree, which would ultimately lead to a less-informative prior for the Allee threshold parameter, $a$, when modelling the Klein et al., data.
[…]
The mean probability of success is an exponential saturating function of $(\text{{attacks per acre}})_{y}$, parameterized with the Allee threshold, denoted $a$:
$\mu_{y,\text{{allee}}}=1-\exp\left(-2.3\cdot\frac{{(\text{{attacks per acre}})_{y}}}{a}\right).$ (32)
Recall that the Allee effect sub-model is first fit to the Boone et al., data (Fig. 8). The subsequent estimate of $a$ is used to set an informative prior for the outbreak model (which is constituted by sub-models and the integrator model, which is then fit to the Klein et al., data. Specifically, we choose a half-normal prior with a standard deviation equal to half of the posterior mean of $a$.
The normal distribution is unrealistic because it implies that probabilities outside of the interval $[0,1]$ are possible. The normal distribution assumes constant variance, whereas the Boone et al., data clearly shows heteroskedasticity. That being said, a more realistic error structure would unnecessarily complicate the model. The normality of errors and constant variance are unimportant assumptions when trying to estimate the conditional expectation (Gelman and Hill,, 2007, Sec. 3.6), which is exactly what we are trying to do — only the mean probability of success (i.e., Eq. 32) is used to model the Klein et al., data.
A.7 Integrator model
The integrator model combines the deterministic skeletons of all sub-models (Appendices A.3A.6) to predict the next year’s gallery starts, solely from this year’s gallery starts. In doing so, we can estimate several parameters that cannot be estimated from singular sub-models: $a$, $s_{0}$, $\lambda_{0}$, $\lambda_{1}$, and a residual noise parameter, $\sigma_{\text{resid}}$.
First, we predict the distribution of successful beetle attacks among diameter classes. The predicted mass attack preference is
$\hat{\text{preference}_{j,y}}=\begin{cases}\beta_{0}+\beta_{1}\cdot\text{{DBH}}_{j}&\text{if }\text{{DBH}}_{j}>-\beta_{0}/\beta_{1}\\ 0&\text{otherwise}.\end{cases}$ (33)
Note that the $\text{preference}_{j,y}$ has a “hat”, which will be used throughout this section to distinguish model-based predictions from data.
[…]
$\hat{(\text{succ. attacks per acre})_{j,y}}=\frac{(\text{succ. attacks per acre})_{y}\cdot\hat{(\text{preference})_{j,y}}\cdot(\text{prop. area})_{j,y}}{\sum_{k=1}^{J}\hat{(\text{preference})_{k,y}}\cdot(\text{prop. area})_{k,y}},$ (34)
[…]
Successful beetle attacks are converted into killed trees with
$\hat{\text{dead}}_{j,y}=\text{min}\left(\gamma\;\frac{\hat{(\text{successful attacks per acre})}_{j,y}}{(\text{susceptible surface area per tree})_{j,y}},\text{TPA}_{j,y}\right),$ (35)
and killed trees are converted into emerging females with
$\hat{(\text{{exit density}})_{y+1}}=\sum_{j=1}^{J}\begin{cases}\overbrace{\left(\zeta_{0}+\zeta_{1}\cdot\text{DBH}_{j}\right)}^{\text{Emergence density}}(\overbrace{(\text{KTPA})_{j,y}(\text{surface area per tree})_{j}}^{\text{surface area of killed trees}}),&\text{if }\text{DBH}_{j}>-\frac{\zeta_{0}}{\zeta_{1}}\\ 0,&\text{otherwise}.\end{cases}$ (36)
To predict the number of attacking beetles, we multiply the exit hole density by the predicted dispersal-phase survival. Contrary to Eq. 4 in the simulation model, there is never any immigration:
$\hat{(\text{{attacks per acre}})_{y+1}}=\frac{s_{0}\;\hat{(\text{{exit density}})_{y+1}}}{1+\exp\left(-\lambda_{1}\left((\text{relative surface area})_{y+1}-\lambda_{0}\right)\right)}.$ (37)
To complete the virtual life-cycle and predict the number of successful attacks next year, denoted $\mu_{y,\text{full}}$, the Allee effect is applied to the predicted attack density
$\mu_{y+1,\text{full}}=\hat{(\text{{attacks per acre}})_{y+1}}\times\left(1-\exp\left(-2.3\cdot\frac{{\hat{(\text{{attacks per acre}})_{y+1}}}}{a}\right)\right)$ (38)
Finally, the likelihood of the integrator model is given by a normal distribution with a residual noise parameter.
[…]
Table A.1: Statistical model notation
Symbol Description Units submodel(s)
Subscript
$t$
Tree age in years
N/A
Lodgepole, A.2
j
The DBH-class index, uniquely identifies a tree diameter class
N/A
All
y
The year index,
N/A
All
tree
Pertaining to the lodgepole pine dyanamics sub-model
N/A
Lodgepole, A.2
pref
Pertaining to the size-dependent attack preference sub-model
N/A
Attack pref., A.3
dead
Pertaining to the tree to killed tree sub-model
N/A
Killed trees, A.4
exit
Pertaining to the size-dependent fecundity sub-model
N/A
Fecundity, A.5
allee
Pertaining to the Allee effect sub-model
N/A
Allee effect, A.6
Data
[…]
Upper age limit of trees within a particular diameter class
years
Lodgepole, A.2
[…]
Lower age limit of trees within a particular diameter class
years
Lodgepole, A.2
$\text{TPA}_{j}$
Density of trees within the ages corresponding to the $j^{th}$ diameter class
[…]
Lodgepole, A.2
$(\text{{preference}})_{j,y}$
Mass attack preference, defined as the attack density within a particular diameter class, divided by the attack density across all trees (within a year)
unitless
Attack pref., A.3
$\text{{DBH}}_{j}$
Diameter at breast height
inches
Attack pref., A.3
$(\text{{KTPA}})_{j,y}$
Killed trees per acre
[…]
Killed trees, A.4
[…]
Gallery starts per acre
[…]
Killed trees, A.4
$(\text{surface area per tree})_{j}$
Susceptible surface area per individual tree, defined by Eq. 1
$\text{meters}^{2}\cdot$ $\text{trees}^{-1}$
Killed trees, A.4
$(\text{{exit density}})_{y}$
Emergence holes per acre
[…]
Fecundity, A.5
$J$
Total number of unique diameter classes,
unitless
Fecundity, A.5
$Y$
Total number of years in the data set
unitless
Fecundity, A.5
$(\text{{attacks per acre}})_{j,y}$
Attacking female beetles per acre
[…]
Allee effect, A.6
[…]
Proportion of total susceptible tree surface area belonging to a particular diameter class
unitless
Integrator, A.7
Derived quantities
$\mu_{j,\text{tree}}$
Expected tree density in the $j^{th}$diameter class
[…]
Lodgepole, A.2
$\mu_{j,y,\text{{pref}}}$
Predicted mean of the mass attack preference
unitless
Attack pref., A.3
$\sigma_{j,y,\text{{pref}}}$
Standard deviation of the mass attack preference
unitless
Attack pref., A.3
$\mu_{j,y,\text{{dead}}}$
Predicted mean of the killed trees density
[…]
Killed trees, A.4
$\sigma_{j,y,\text{{dead}}}$
Standard deviation of the killed trees density
[…]
Killed trees, A.4
$\mu_{y,\text{{exit}}}$
Predicted mean of emergence hole density
(exit $(\text{exit holes})\text{acres}^{-1}$
Fecundity, A.5
$\mu_{j,y,\text{{allee}}}$
Mean proportion of successful attacks
unitless
Allee effect, A.6
$s_{y}$
Actual survival probability for beetles during the dispersal phase
unitless
Integrator, A.7
$\lambda_{y}$
Stand-dependent dispersal factor; actual mortality increases as the proportion of remaining susceptible surface area decreases
unitless
Integrator, A.7
$\mu_{y,\text{full}}$
Expected density of successful beetle attacks (across all DBH classes)
[…]
Integrator, A.7
Model parameters
$K$
Tree carrying capacity
[…]
Lodgepole, A.2
$m$
Annual probability of tree death (not caused by MPB)
unitless
Lodgepole, A.2
$\sigma_{\text{tree}}$
Noise parameter
[…]
Lodgepole, A.2
$\beta_{0}$
Y-intercept of the piecewise linear preference-DBH function
unitless
Attack pref., A.3
$\beta_{1}$
Slope of the piecewise linear preference-DBH function; the DBH threshold for attacks is $-\beta_{0}/\beta_{1}$
unitless
Attack pref., A.3
$\sigma_{\text{{pref0}}}$
noise parameter; accounts for a baseline level of variability in mass attack preference
unitless
Attack pref., A.3
$\sigma_{\text{{pref1}}}$
Noise parameter; accounts for increasing variability with smaller sample sizes in mass attack preference
unitless
Attack pref., A.3
$\gamma$
Proportionality constant for number of killed trees; the reciprocal of the average attack density
$\text{meters}^{2}\cdot$ $(\text{gallery starts})^{-1}$
Killed trees, A.4
$\sigma_{\text{{dead0}}}$
noise parameter; the effect of the relative attack density on the standard deviation
$\text{meters}^{2}\cdot$ $(\text{gallery starts})^{-1}$
Killed trees, A.4
$\zeta_{0}$
Y-intercept of the piecewise linear emergence-DBH function
[…]
Fecundity, A.5
$\zeta_{1}$
Slope of the diameter-emergence regression; the DBH threshold for emergence is $-\zeta_{0}/\zeta_{1}$
[…]
Fecundity, A.5
$\sigma_{\text{{exit}}}$
Standard deviation parameter for emergence holes
[…]
Fecundity, A.5
$a$
Allee threshold (density of beetles at which probability of attack success is 0.9)
[…]
Allee effect, A.6
$\sigma_{\text{{allee}}}$
Standard deviation parameter for proportion of successful attacks
unitless
Allee effect, A.6
$s_{0}$
Maximum survival probability for beetles during the dispersal phase
unitless
Integrator, A.7
$\lambda_{1}$
Sensitivity of dispersal mortality on tree surface area; controls the speed of the transition from low to high dispersal mortality as trees are killed.
unitless
Integrator, A.7
$\lambda_{0}$
Transition point parameter; controls the level of tree death at which dispersal mortality becomes significant
unitless
Integrator, A.7
$\sigma_{\text{resid}}$
Residual noise parameter, to account for error in the predicted density of successful beetle attacks next year
[…]
Integrator, A.7
Appendix B Additional model justification
B.1 Mass attack preference does not change
One potential shortcoming of our model is that preference only depends on tree diameter, whereas previous papers have shown that MPB’s preference for large-diameter trees increases as outbreaks progress (Boone et al.,, 2011; Howe et al.,, 2022). To investigate further, we separately fit our preference sub-model to each year of data and plotted $\beta_{1}$ (the slope of the preference $\sim$ DBH regression) against beetle pressure, as measured by the density of emerging females, and the number of emerging females per unit of susceptible tree surface area (Fig. B.1). There was no clear pattern, indicating that dynamic changes to the DBH preference are small or nonexistent. Similarly, there was no relationship between beetle pressure and the residuals of our original preference model (Fig. B.2).
[…]
Figure B.2: The mean and 95% credible intervals of the $(\text{preference})_{j,y}$ residuals — from the size-dependent preference sub-model — as a function of beetle pressure. Two metrics of beetle pressure are used: Beetles, i.e., recently emerging female beetles (units: thousands per acre), and Beetles/host, i.e., the quotient of emerging female density and within-DBH-class surface area (units: females per acre per square meter). There is no clear relationship between residual values and beetle pressure. Data from Klein et al., (1978).
There are a couple of ways to explain the absence of evidence for dynamic DBH-preferences. First, it could be that MPB’s large-tree preference “kicks in” at an early, unobserved stage of the Klein et al., outbreak. In Fig. 6. of the publication by Boone et al., (2011), the ratio of DBH of trees entered and not entered increases from 1.2 to 1.6 as an outbreak progresses from endemic to eruptive conditions. However, based on Boone et al.,’s classification scheme, the Klein et al., outbreak is approximately in the eruptive condition until the very end of the outbreak (calculations in supplementary files). Second, dynamic changes to DBH-preferences could be small. In Fig. 3D of Howe et al., (2022), the mean DBH of attacked trees only increased a few centimeters as infestation severity increased from 1 to 1000 attacked trees per hectare.
B.2 Increasing attack density does not affect population dynamics
Figure 11 in the main text shows that the density of attacks on killed trees (e.g., gallery starts per square meter) increases over time. This pattern can potentially be explained by a positive relationship between attack density and beetle population size, an assumption made by several models in the literature. To investigate further, we fit the killed-tree sub-model (Appendix A.4) to each year of data and plotted $1/\gamma$ (the model-based estimate of attack density) against beetle pressure, as measured by the density of emerging females, and the number of emerging females per unit of susceptible tree surface area (Fig. B.3). The plots are suggestive of a positive relationship if the data from 1971 are excluded, but the nature of the relationship is ultimately inconclusive.
Even if beetle pressure does affect attack density, this has negligible influence on population dynamics. The killed-tree sub-model provides an excellent fit to the data, and the residual variation cannot be explained by beetle pressure (Fig. B.4). This is likely because variation in attack density is swamped by variation in $\frac{(\text{succ. attacks per acre})_{j,y}}{(\text{surface area per tree})_{j,y}}$ (see Eq. 8 in the main text).
Figure B.3: The mean and 95% credible intervals of $1/\gamma$ — the idealized attack density in killed-tree sub-model — as a function of beetle pressure. To obtain multiple estimates of $\gamma$, the killed-tree sub-model was fit independently to each year of data. Two metrics of beetle pressure are used: Beetles, i.e., recently emerging female beetles (units: thousands per acre), and Beetles/host, i.e., the quotient of emerging female density and within-DBH-class surface area (units: females per acre per square meter). Data from Klein et al., (1978).
Figure B.4: Mean and 95% credible intervals of $(\text{KTPA})_{j,y}$ residuals from the killed-tree sub-model, as a function of beetle pressure. Two metrics of beetle pressure are used: Beetles, i.e., recently emerging female beetles (units: thousands per acre), and Beetles/host, i.e., the quotient of emerging female density and within-DBH-class surface area (units: females per acre per square meter). Data from Klein et al., (1978).
B.3 The Allee effect depends only on beetle density
Our model states that the success probability (i.e., the probability that an attacked tree is successfully killed) only depends on beetle density. Several previous papers have modelled the success probability as an increasing function of the quotient of beetles and susceptible trees, a simplifying assumption made for mathematical tractability. In this appendix, we argue that these models 1) are unrealistic, and 2) themselves show that the quotient of beetles and trees is superfluous for modelling the Allee effect, as a practical matter.
A good example of a model with non-separable beetles and trees is provided by Goodsman et al., (2017). In this model, the densities of beetles and susceptible trees are denoted $B$ and $S$, respectively. In each generation, only an $\alpha$ proportion of susceptible trees are attacked. As the authors write, “In reality, because of the clustered nature of attacks in space, most trees in a stand are not attacked and attacks are focused on a small subset of trees (Logan et al.,, 1998) some of which may receive many more attacks than others.” The probability that an attacked tree is killed is an increasing function of $m=B/S$.
Each attacked tree receives a number of beetles that is drawn from a negative binomial distribution. The distribution is parameterized with the mean number of beetles, $m=B/(\alpha S)$, and a dispersion parameter $k$ (note that we are redefining some parameters from the main text). Both parameters modulate aggregation — small $k$ and small $\alpha$ both correspond to high aggregation, where a small number of trees receive a large number of beetles. Finally, a tree is killed if the number of beetles it receives is greater than the threshold parameter $\phi$. With these dynamics, the probability of success (i.e., the probability that an attacked tree is killed) is the probability mass across all attacks with more than $\phi$ beetles:
$F(m;\phi,k)=\sum_{i=\phi+1}^{\infty}NB(i;m,k).$ (40)
This is the complementary CDF of the negative binomial distribution.
While this model may be attractive from a mathematical point of view, it has several unrealistic assumptions: the proportion of attacked trees, $\alpha$, does not scale with the number of beetles; the per-surface-area attack density may be arbitrarily large, when in reality it is constrained by gallery defense and anti-aggregation pheromones; and brood size does not increase once the attack threshold $\phi$ is met (see Raffa and Berryman,, 1983, Fig. 6 for counter-evidence). More seriously, the model leads to unrealistic population dynamics. If aggregation is too strong (i.e. $k\ll 1$), then the beetles will always be able to find the few fractional trees that slip into the susceptible age class each year; this leads to a large stochastic equilibrium of beetles that is much larger than endemic population seen in nature. The solution, it would seem, is to accept that aggregation is not so strong. But weak aggregation makes it difficult for beetles to invade a mature forest; $m=B/S$ is small, and the beetles are spread evenly across trees, so the threshold $\phi$ is never exceeded. Counter-intuitively, the beetles get worse at invading as a forest becomes denser since $m=B/S$ becomes smaller. In our experience, with realistic values of $\phi$ and $\alpha$, it is difficult to find values of $k$ where beetles experience realistic boom-bust dynamics — where beetle populations will be able to invade, but will later recede to an endemic state.
While acknowledging certain limitations in the Goodsman et al., (2017) model, we will accept it as a premise to illustrate that the probability of mass attack success is predominantly influenced by beetle density. Our Allee effect sub-model with a single parameter (Appendix A.6) is well-approximated by the comparatively parameter-rich model of Goodsman et al., (2016), when the latter model’s parameters are given realistic values.
We examine the relationship between the success probability, here given by Eq. 40, and our three state variables: $B$, $S$, and $m=B/S$. More specifically, we examine how the nature of this relationship changes as the parameters $\alpha$ and $k$ change. Figure B.5 shows that the probability of success only shows a weak dependence on $S$ when the dispersion parameter is large (i.e. $k>1$) and the attacked fraction of trees is small (i.e. $\alpha\ll 1$). This is precisely the parameter combination that best corresponds to reality (Table B.1; Fig. B.6): beetles spread themselves evenly (i.e. $k>1$) across a small fraction of susceptible trees (i.e. $\alpha\ll 1$), which they can successfully kill. Success probability remains high unless $S$ is unrealistically large.
Figure B.5: The probability that an attack tree is killed, calculated via Goodsman et al.,’s 2017 model (Eq. 40), as a function of $B$ (female beetles per acre), $S$ (susceptible trees per acre), and $m=B/S$. The range for beetle density was selected based on the Klein et al., dataset, and the range for susceptible tree density was selected based on the literature: the density of lodgepole pine forests with susceptible trees rarely exceeds 1500 trees/ha, or equivalently, $\approx 600$ trees/acre (Nigh et al.,, 2008; Pfeifer et al.,, 2011; Meddens et al.,, 2011). For panels A and B, the other state variable is fixed at $S=500$ and $B=3000$ respectively. To reduce the parameter space that we must numerically explore, we fix $\phi=176$, which is the attack density threshold given by Raffa and Berryman, (1983) — approximately 40 females per square meter —- times the susceptible surface area of a 10-inch tree — approximately 4.4 square meters.
Figure B.6: Estimates of $\alpha$, the proportion of susceptible trees that are attacked. The data are sourced from Boone et al., (2011, Table 1). The median of is $0.009$.
Source $\overline{\mu}$ $\mu_{\text{2.5\%}}$ $\mu_{\text{97.5\%}}$ $\overline{k}$ $k_{\text{2.5\%}}$ $k_{\text{97.5\%}}$
Peterman, 1974, Thesis, Fig. 4, Plot: Elk 119.6 104.6 136.6 9.7 5.2 16.1
Peterman, 1974, Thesis, Fig. 4, Plot: Lean 84.8 75.3 95.6 3.9 2.8 5.2
Peterman, 1974, Thesis, Fig. 4, Plot: Pars 75.1 63.3 89.1 2.9 1.9 4.1
Peterman, 1974, Thesis, Fig. 4, Plot: Terr 76.5 66.5 87.7 3.6 2.5 5.1
Safranyik, 1968, Thesis, Fig. 32, Plot: Elk creek 32.1 23.3 43.9 6.5 1.7 16.9
Safranyik, 1968, Thesis, Fig. 32, Plot: Horsethief creek 23.8 20.9 26.9 23.8 7.3 68.8
Shepard, 1965, Can. Ento., Fig. 4 46.4 37.4 57.7 3.2 1.7 5.3
Waring and Pitman, 1985, Ecol. Fig. 1 92.6 78.6 110.2 2.1 1.5 2.9
Table B.1: Mean and dispersion parameters ($\mu$ and $k$ respectively) for the negative binomial distribution, fit to the distribution of attack density (units: females per square meter) across successfully and unsuccessfully attacked trees. Posterior means and 95% credible intervals endpoints are displayed. The data are sourced from Peterman, (1974); Safranyik, (1968); Shepard, (1965); Waring and Pitman, (1985). The dispersion parameter $k$ may be scale-dependent, so it is important to note the sizes of the study areas. All studies were conducted at the stand-level, but the sizes of the study areas are not directly comparable due to disparate sampling methodologies (i.e., transect vs hierarchical spatial sampling).
Appendix C Model fitting details and diagnostics
We checked a standard suite of summary statistics to provide evidence that the Markov Chains had converged to the global posterior and were efficiently sampling. In particular, $R<1.1$ for all parameters (i.e., all chains had mixed), the effective sample size per iteration was greater than 0.001 (i.e., efficient sampling), the energy Bayesian fraction of missing information (E-BFMI) was less than 0.2 (i.e., the model was not severely misspecified), and the proportion of divergent trajectories was far less than 1% (i.e., the posterior geometry did not lead do biased estimation). This diagnostic analysis is provided in the R markdown file model_diagnostics.Rmd and its output model_mcmc_diagonstic_printoff.txt, in the supplementary files. See Gelman et al., (2014, Ch. 6) for more information on the aforementioned summary statistics.
To show that the marginal prior distributions had negligible influence on the posterior distribution, we computed the posterior contraction.
[…]
a parameter-specific relative influence of the data on parameter estimation. For nearly every parameter in every model, the prior contraction was above 0.9, implying negligible prior influence. The exception is the Allee threshold parameter ($a$) in the complex outbreak dynamics threshold. The posterior contraction analysis is provided in the R markdown file model_diagnostics.Rmd and its output model_post_contract_printoff.txt, in the supplementary files.
Parameter Short description Mean SD $\text{CI}_{2.5\%}$ $\text{CI}_{97.5\%}$
[…]
$\sigma_{\text{pref0}}$ pref. baseline noise 0.630 0.0822 0.487 0.810
$\sigma_{\text{pref1}}$ pref. sampling noise 0.152 0.113 0.00478 0.412
[…]
$\sigma_{\text{dead}}$ dead tree noise 0.00508 0.000668 0.00399 0.00660
[…]
$\sigma_{\text{exit}}$ emerge noise 3.61 1.54 1.79 7.47
$s_{0}$ max dispersal survival 0.474 0.0676 0.363 0.627
$\sigma_{0}$ residual noise 365 89.7 224 571
$\lambda_{\text{resid}}$ dispersal threshold 0.652 0.0339 0.582 0.705
$\lambda_{1}$ dispersal sensitivity 49.7 96.6 17.2 262
a Allee threshold 392 331 12.1 1211.68175
K forest carrying capacity 904 313 413 1630.37825
m annual tree mortality 0.0518 0.00637 0.0376 0.0634
$\sigma_{\text{tree}}$ tree abund. noise 0.482 0.211 0.240 1.02
Derived quantity Short description Mean SD $\text{CI}_{2.5\%}$ $\text{CI}_{97.5\%}$
$\sigma_{\text{tree}}$ tree abund. noise 0.482 0.211 0.240 1.02
$-\beta_{0}/\beta_{1}$ susc. DBH threshold 7.43 0.643 6.04 8.43
$1/\gamma$ attack density 66.2 3.78 59.2 74.1
$\zeta_{0}+\zeta_{1}*10$ emerge density (10 in. DBH) 244 25.4 192 297
Table C.1: Summary statistics for all model parameters, including the Mean, standard deviation, and parameter credible intervals ($\text{CI}_{2.5\%}$, $\text{CI}_{97.5\%}$). Here, the parameters come from the full hierarchical model (Appendix A) fit to the Klein et al., dataset. The table also includes derived quantities, including the DBH of the smallest susceptible trees, the model-estimated attack density (gallery starts per square meter), and the emergence density (females per square meter) on a tree with a 10-inch DBH, which is approximately the average DBH of attacked trees in the Klein et al., dataset. Finally, the table contains fixed parameters, whose values were obtained via literature search. For parameter units, and the summary statistics of sub-model noise parameters, see supplementary Table A.1.
Parameter Short description Mean SD $\text{CI}_{2.5\%}$ $\text{CI}_{97.5\%}$
a Allee threshold 2492.2411175 569 1631.332 3809.797
$\sigma_{\text{Allee}}$ Pr. success noise 0.207 0.0300 0.158 0.275
Table C.2: Summary statistics for Allee effect model parameters, including the Mean, standard deviation, and parameter credible intervals ($\text{CI}_{2.5\%}$, $\text{CI}_{97.5\%}$). Here, the parameters come from the Allee effect sub-model (Appendix A.6) fit to the Boone et al., dataset.
Appendix D Supplementary Figures
Figure D.1: Approximate dispersal mortality across time in the Klein et al., data. The dispersal mortality was estimated as the quotient of gallery starts and preceding emergence holes. The approximation will be better when emergence density is larger, because the Allee effect will be less important, and local population dynamics — compared to dispersal — will be more important.
Figure D.2: Population maps provide evidence for compensatory density dependence at the tree level. The data comes from Alberta’s disk surveys (described by author N/A,, 2020, and provided by Goodsman et al.,, 2018). The estimate for attack density at time $t+1$ involves the constants $(2/3)$ (the average proportion of the brood that will be female) and $0.3$ (an estimate of dispersal-phase mortality from the Klein et al., (1978) dataset, see Fig. D.1).
Figure D.3: Sample time series of MPB-infested trees from British Columbia’s Aerial Overview Surveys. Each panel is a randomly selected $\text{1 km}^{2}$ patch from a subset of locations where cumulative infestations were in the upper quartile. To estimate the number of infested trees, we multiply the proportion of infested stems by Live aboveground biomass (units: $\frac{\text{Live aboveground biomass (units: Mg/ha)}}{10}$, which is an approximation for pine stem density that was validated by Koch et al., (2021, Appendix A), using the data and analyses of Nigh et al., (2008). The biomass data is sourced from Beaudoin et al., (2014).
Figure D.4: In 1970, emergence density was high but subsequent attack density was low. Anomalous weather conditions during the flight period could be responsible for this pattern, but the relevant weather variables were unexceptional in 1970 (the bold, teal line). The maximum summer temperature was consistently greater than 18.3°C, the lower limit for beetle flight (McCambridge,, 1971). The estimated wind speed often exceeded 8 km/hour, the proposed upper threshold for beetle flight (Gray et al.,, 1972); however, the wind speeds in 1970 were not anomalous in comparison to the other outbreak years. Weather data was obtained from the Biosim program (Régnière et al.,, 2014).
Figure D.5: In 1970, emergence density was high but subsequent attack density was low. The decline could theoretically be attributed to poor environmental conditions that weakened the beetles, despite not being lethal (because they still emerged). However, the developmental period leading up to 1970 had unexceptional weather conditions. The minimum winter temperature is related to larval mortality (here, sub-lethal damage), and degree days are related to emergence and univoltinism. The minimum winter temperature is approximately -30°C, which is substantially greater than the proposed threshold of -40°C (Safranyik,, 1998) for significant overwintering mortality. There are approximately 1250 degree days from Aug. 1 to Jul. 31, far beyond the proposed threshold of 833 for the population to be univoltine (Reid,, 1962). Weather data was obtained from the Biosim program (Régnière et al.,, 2014).
Figure D.6: The stable epidemic equilibrium (i.e., the intersection of the nullclines with the highest beetle density) has a very small basin of attraction. All other initial conditions lead to the endemic equilibrium.